by Champlin Champlin

Courthouse Lobby Security Camera Install

11/12/19 – SOSS approved for Camera Install Project
12/09/19 – Champlin met with Matthew Butz at the Courthouse to discuss scope of work.
12/17/19 – Champlin Contacted Beth Johnson with the Historical Conservation Office and approval was not needed for this project.
01/30/20 – Project okay to archive, the work is complete.

by Champlin Champlin

CH – 6th Floor Office Build Out

08/27/19 – County requests onsite meeting to review scope.
09/10/19 – Champlin received approved SOSS & PO.
09/17/19 – Champlin & ThermalTech site visit.
09/27/19 – Check set given to Hamilton County for approval.
10/03/19 – CDs set given to Hamilton County and issued for permit.
12/03/19 – Bulletin #01 sent to County and Triton Services.

12/31/19 – Door and hardware submittals approved and sent to Triton

01/02/20 – CO #4 signed and sent to County

02/03/20 – Champlin performed Punchlist visit

02/04/20 – Punchlist report sent to County and Triton

02/14/20 – Champlin performed Back Punch visit

02/21/20 – Backpunch report sent to County and Triton

05/15/20 – Pay app signed and returned to Triton

by MCB MCB

CH – West Entry Security Screening Renovations

04/24/19 – HCFD requests SOSS from TEI for the project.  TEI submits SOSS and receives signed SOSS and PO.  TEI requests kickoff meeting the HCFD and all parties .  Project start pending project kickoff meeting.

05/29/19 – Design progressing and TEI anticipates that 100% documents will be issued to HamCo no later than mid week (week of June 3).

06/26/19 – Drawing submitted to HamCo for review and no comments received.  TEI submits to IBI for plan review and permit on 06/28/19.

07/24/19 – TEI receives plan review approval from IBI, and secures the electrical permit for trades to work on the project.  Core drills complete and trades working on electrical modifications for the project.  Hamilton County trades to notify ThermalTech when they are ready for the rough in inspection.

08/27/19 – work completed , inspections passed, and the security scanners have been relocated.

by TTC TTC

CH – Room #160 HVAC Study

07/16/18 – Sent Tony and updated summary of the fieldwork and load calculations associated with Room #160.  Provided basically four options for possible HVAC changes that may help with cooling issues.

  • Unit appears to size large enough for the design day load
  • Add reheat to all interior zones for additional comfort control
  • Replace the unit and add zoning
  • Add DX unit to just #160
  • Relocate Judges to rooms that match their cooling request

08/01/18 – AAB provides hard quote for work based on ThermalTech drawings; they are estimating three to four days of field work to rebalance everything and test it.
08/03/18 – Tony contacts ThermalTech about moving forward with project.
08/23/18 – TTC sent a follow up email to Tony and Rob outlining what Honeywell found while on site performing additional research.  Al the existing VAV boxes are pneumatic and any control strategy with intelligent reset design is not possible unless they are upgraded to DDC boxes.  Hot water not available on this floor but is available in the basement with the unit.  TTE provided a general estimate to upgrade the boxes and extend hot water reheat to the new spaces.  What we have not found is what temperature the hot water is disabled based on OAT and what that may do to this design.
09/27/18 – Court decided to wait for funding before they do anything and what they have now will work for the foreseeable future.
10/23/18 – This project archived by TTC.

by GBBN GBBN

CH – Bronze Door Cleaning

06/21/2018 – Tony Matre contacts GBBN to inquire about this past project and the status of the original developmental drawings.  GBBN indicates all documents are still in possession of GBBN.  Tony would like to Bid  the Advertisement no later than November 16, 2018.

by GBBN GBBN

CH – Room #205 Renovation

04/02/18 – Facilities notifies GBBN that Municipal Courts has approved the project, estimated with Project Request #1949, for submission of an SOSS, to be performed by Trades. SOSS to cover the scope of the Project Request #1949.  GBBN solicits input from TEI.  Facilities confirms Trades to handle project including systems furniture scope.
04/03/18 – TEI sends input for SOSS to GBBN.
04/04/18 – SOSS issued to Facilities.
05/29/18 – TTC and SJ discussed starting this project after the JC Recovery is finished. The SOSS is in place for our firms. GBBN will develop the floor plan and TEI will star the load calcs and preliminary work right after the floor plans are ready.
04/03/18 – TEI sends SOSS input to GBBN.
04/03/18 – GBBN submits SOSS to Facilities.
05/10/18 – GBBN receives approved SOSS and Purchase Order from Facilities.
06/26/18 – TEI and GBBN discussed schedule for this project to be completed the first couple of weeks of July 2018. TTC, Coty and Barth scheduled to work on this project.
07/25/18 – GBBN has construction document set started.  GBBN questions were answered from Facilities to keep documents moving ahead. TTE checking load and ventilation calcs.
08/20/18 – TEI confirms that no fire protection modifications are required from the existing layout and configuration.
08/28/18 – Facilities requests three minor revisions to Owner Review set of documents after review by Municipal Court Administrator.
08/29/18 – TEI and GBBN have submitted final drawings for Owner Review.  Facilities confirms only one small revision, of adding data wall mounted outlet for future printer/copier, is required before issuance for Permit.  GBBN and TEI agreed upon 09/04/18- Issuance for Permit date in issue block.
09/05/18 – GBBN and TEI submit Applications for Building and Mechanical Permits to the City of Cincinnati.
09/26/18 – TEI receives code comments and re-submits drawings to IBI.
10/03/18 – GBBN receives letter from the City of Cincinnati requesting additional information for permitting.  This information is far in excess of any required by the City for comparable projects since GBBN has been doing Facilities projects.
10/17/18 – GBBN requests proposal from The Preview Group to provide part of the information requested by the City mentioned above.
12/21/18 – GBBN submits City of Cincinnati Permit Response Letter.
01/07/19 – GBBN pays for and obtains Building Permit from City of Cincinnati. Facilities agrees reimbursable expense to be billed to General Services SOSS.
01/29/19 – ThermalTech obtained the permit and delivered to John Nester for trades to do the work.
02/26/19 – Work is in progress with the county trades.  As of 2/15/19, there was no electrical roughed in to date.  Wall framing was nearing completion.  TEI tells John Nester to let us know when the trades are ready for a rough in inspection with IBI and TEI will schedule the inspection.
03/08/19 – GBBN performs site visit with Facilites. Inspector questioned location of two sprinkler heads. TEI responds to question with data input in email to Facilities.
03/26/19 – Work continues to progress in area.  Electrical rough in inspection was passed and work continues in space.  Trades (JN) to notify TEI when to call for final on the project.
04/24/19 – HamCo to notify TEI when to call IBI for final on the project.
05/29/19 – TEI was notified to contact IBI for final electrical inspection.  IBI scheduled for final inspection on 5/31.
06/26/19 – Project final IBI approval received.  on 06/25/19, while on site for a different project, the rooms were being moved into the space.
06/27/19 – Project is complete. Ready to archive. Partners can send in CAFM updates.

by THP THP

CH – Critical Façade Repairs (THP F103)

12/28/17 – THP publish SOSS to Hamilton County as requested.
03/16/18 – Hamilton County approved SOSS.
04/26/18 – THP provided owner review documents to the County.
05/24/18 – Revised Owner review documents sent to County.
05/25/18 – County provided comments.
05/29/18 – Documents issued to purchasing for review.
05/30/18 – Purchasing provided comments to the review documents. Project issued ITB 044-18.
06/1/18 – Issued bid documents to County and ARC for publication.
06/5/18 – Documents available to bidders
06/11/18 – Prebid meeting occurred at Courthouse.
06/12/18 – Addendum No.1 and No.2 issued.
06/26/18 – Bid opening occurred with four contractors submitting proposals. Two of the four bids under published estimate. American Façade Restoration is the apparent low bidder.
06/27/18 – THP had phone conversation with AFR to review their bid.
07/02/18 – THP provided a bid summary letter and draft AIA contract for review by the County.
07/05/18 – County published updated AIA contract and documents for review by AFR.
7/25/18 – County published executed AIA contract.
08/21/18 – Background checks approved and submittals in process.
08/30/18 – Preconstruction meeting scheduled for 9/1/18.
09/1/18 – Preconstruction meeting completed.
09/26/18 – American Scaffolding erected swing stage equipment for courtyard repairs.
10/12/18 – Change Order #01 Executed. During the dentil repairs it was discovered the cornice steel supports were heavily corroded. Change order #1 provide new support detail, that is integrated into the dentil repair.
10/25/18 – AFR working on courtyard dentil repairs.
11/16/18 – Change Order #02 Executed. Work to address loose stone, concrete, masonry materials observed during city façade assessments.
11/29/18 – AFR completed Courthouse base bid repairs. Roughly 60% of knock down survey and stone stabliziation work complete. Work scheduled to be completed December 8th (weather permitting).
12/07/18 – AFR completed all knock down work for Change order #02
12/31/18 – AFR completed final window repair for Change order #02.
01/17/19 – THP submitted the certified AFR pay application to the County.
01/18/19 – AFR submitted close out documentation to THP for review.
03/21/19 – THP forwarded final close out documents to the County.
04/25/19 – Project complete

by TEI TEI

COURTHOUSE 6th Floor 2 Hard Shell Offices

07/12/17 – Kickoff Meeting to discuss scope of work.  Measurements and photographs were taken.
07/13/17 – GBBN discusses scope and sends sketch of proposed layout to TEI for SOSS input.
07/17/17 – Facilities confirms this project would require a Walk-Through-Permit application.
08/10/17 – GBBN authors and delivers SOSS to Facilities.  Project is on hold.

by GBBN GBBN

CH – 7th Floor Core and Shell Study

07/17/17 – Kick off Meeting with GBBN and Facilities to walk the space and adjacent areas.
07/19/17 – Facilities confirms extending an elevator to this level would be required as part of the scope.
07/27/17 – Scope of work for this is to just do the study and construction estimate.  Include utilities, elevator, windows, core and shell upgrades.  Can an elevator be installed that just goes between 6 and 7 – is that cheaper.  SOSS requested in less than two weeks, provide scope, cost, investigation, code review and and quick SD type design.
08/08/17 – GBBN  meets with Tim Davis of Aluminum Resources to discuss window replacements and possible criminal justice type window replacements.
08/11/17 – THP issued SOSS for structural support services for renovation study.
08/12/17 – GBBN authors and delivers SOSS to Facilities.
06/27/19 – Ok to archive.

by TEI TEI

COURTHOUSE Roof Y and Z Leak Investigation

Partner Lead: THP


05/05/17 – Hamilton County requested SOSS to investigate roof leaks on the Y and Z roofs along the west elevation.

05/25/17 – THP developing SOSS.

06/08/17 – THP issued SOSS for review.

06/26/17 – County approved SOSS.

07/27/17 – THP plans to schedule work second week of August.

08/25/17 – THP visited site to do visual survey of roofs Y & Z, and interior of 6t and 7th floor spaces below roofs.

08/25/17 – THP visited site to do visual survey of roofs Y & Z, and interior of 6th and 7th floor spaces below roofs.

09/28/17 – Based on weather forecast, THP planning on conducting thermal scans October 3rd.

10/03/17 – THP documented roofs Y & Z using infrared camera.

10/06/17 – THP met Underground Detective to inspect all drains at roofs Y & Z.

10/25/17 – Roof cuts scheduled, weather permitting.

10/26/17 – THP visited site to perform investigative roof cuts.

11/28/17 – THP finalized draft of report summarizing findings and recommendations.

by GBBN GBBN

CH – Law Library Renovation

05/15/17 – Hamilton County Facilities informs GBBN of pending approval for project. Facilities requests SOSS.
05/16/17 – GBBN contacts consultants about project and clarifies scope of work included.
05/19/17 – GBBN contacts former Law Librarian to discuss intents in 2016 verses any changes due to change in Librarian.
05/22/17 – GBBN performs site visitation and discussions with Lauren Morrison, new Librarian.
05/24/17 – GBBN and TEI visit site, Facilities to discuss existing drawings, and the Mechanical Space to the north of the existing Stack Area.
06/28/17 – GBBN, IPM, TEI and Facilities hold kick-off Meeting with Law Librarian for Kick-off Meeting.
07/05/17 – GBBN holds in-house discussion to finalize scope and design issues such as Alternate Bids.
07/19/17 – ThermalTech discussed scope change on project with GBBN. More work and revised SOSS required.
08/07/17 – Meeting with Facilities and Law Librarian to discuss additional scope. GBBN asked to author additional SOSS with separate F-number.
08/14/17 – GBBN authors SOSS, for extended scope, and delivers to Facilities.
08/21/17 – GBBN met with Paul Hauser, Facilities Trades to discuss plaster repair scope and material recommendations.
08/29/17 – GBBN sends background drawings and Title Block to TEI.
08/30/17 – GBBN and TEI met regarding MEP coordination issues. Money approved for this project yesterday.
08/31/17 – Discussed scope on project. BW said to include two RR for upgrade (could be alternate), access ramp, move computer door, update break room, expand conference room, update browsing area, investigate covered windows (check on this ), potentially add speaker and microphone if room is enlarged, up to TEI to choose VAV height for this unit.
09/12/17 – GBBN receives approved SOSS for extended scope of services.
09/15/17 – GBBN has internal meeting to discuss full scope now that second SOSS for extended services has been approved.
09/21/17 – GBBN forwards lighting cutsheets and specifications for proposed lighting fixture selections to TEI.
09/25/17 – ThermalTech HVAC schedule for internal review tomorrow on this project. The basic load calcs are complete. TTC to follow up on plumbing and fire protection design.
10/17/17 – GBBN and Facilities meet with Law Librarian for Design Progress Meeting. Final Sketch-Up model forwarded to Librarian for her use in presenting to the Board, and ThermalTech for use in coordination. New window units for the State of Ohio Room will be designated as an Alternate Bid.
10/27/17 – Law Librarian forwarded information concerning the microfiche cabinet size and numbers.
11/09/17 – GBBN internal coordination meeting. Microfiche locations coordinated and power/data routing proposed to TEI.
11/10/17 – GBBN and TEI meet with Leesman Lighting to discuss light fixture selections and specifications.
11/16/17 – Law Library Design Review Meeting at Courthouse. Microfiche layouts coordinated.
11/30/17 – Final Review set due Dec 14 with Owner and Tenant.
03/14/18 – Prebid meeting held for this project at the Courthouse. It was well attended.
03/24/18 – Addendums 1-3 issued for this project. TEI relocated the Ohio Room diffusers for a cleaner install and verified the two ladder locations will work for this job.
03/28/18 – Bid Opening held at CAB, Room #507.  Bids received were in excess of the 10% above Published Estimate and therefore rejected as mandated by Ohio Revised Code.  Facilities to consider next step for reissuing for bidding.

04/25/18 – ThermalTech has started modifying the existing drawings for the new value engineered bid documents.

by TTC TTC

CH – AHU-51, 53, 54 Replacement

01/25/17 – TEI submits SOSS for project. Project pending SOSS approval and PO.
11/17/16 – Hamilton County requested SOSS.
02/16/17 – TEI, Facilities and Building Manager meet on site to kick off project.
02/22/17 – Trane has provided first cut sheet on modular unit for this job, we are trying to fit it into different locations, we are going with a 20 ton unit.
03/01/07 – ThermalTech sent preliminary equipment locations to the County and requested fieldwork to follow up on schematic design. We could potentially have a design meeting at this point and solidify where the unit will be installed.
03/09/17 – ThermalTech updated the floor plans after another field trip down to the site and proposed a final location of one 20-ton unit on the 6th floor with locations for new ductwork and a new wall to enclose the new unit. The county to follow up with Sheriff about using this space for a new unit. TTE moving forward with schedules and Trace runs.
03/15/17 – Revised cost estimates sent to Bert. We are still overbudget but not as badly after electrical modified their cost. We are waiting on revised data from GBBN and THP prior to making additional changes to the estimate. We can make the 6th floor unit work but now its a capital investment issue. Load calcs are complete and we can serve these spaces with a 20 ton, 10,000 cfm unit and we believe we can fit all the adjoining ductwork together to make that happen, the unit would go on the 6th floor.
03/17/17 – TEI contacts GBBN and THP, requesting Architectural and Structural respectively, assistance on this project and included the 5th and 6th floor proposed work drawings for reference.
03/20/17 -50% design dwgs were sent for review to Bert, Tony and Andy. The construction estimate is a little higher than the 192,000 originally budgeted. We are currently trying to finalize the general and structural numbers.
03/20/17 – GBBN sends estimating input to ThermalTech on Architectural implications of the project.
03/21/17 – Spoke to John Milar at THP and we will meet on Thursday to review the AHU locations for the Alternate bids.
03/23/17 – THP – Project explored different approaches not to provide platforms. Meeting now planned to evaluated platforms on 3/23/17.
04/26/16 – Jail Holding unit was added to the design and added as an Alternate to the job. New review drawings were sent to County on 4/25. Additional budgeting was requested for additional design and change orders by owner.
04/27/17 – THP reviewed field conditions and analysis original drawings. Proposed unit and location do not require modification to the structural frame.
05/12/17 – Project is 90% designed but in holding pattern awaiting bid results from the JC Platforms.
06/26/17 – Received revised budgets from Trane on this project and worked on relief air scenario a little bit more. Bid opening could be in July for this project.
06/28/17 – Render the Jury Room for Judge Shannon to see the new ductwork and VAV box.
07/20/17 – GBBN and TEI meet with Facilities for a Field Verification session on site.
07/24/17 – Design team met on site with GBBN to start wall and ceiling work. Some abatement is required at windows we are using for this project. Wall layout confirmed. ThermalTech updated HVAC drawings and they are ready for review. It was decided the Jury room would have an L Soffit instead of a new ceiling, TEI to work with GBBN on this and add access panels for VAV box maintenance.
07/24/17 – GBBN meets with Pinnacle Engineering and Facilities to examine the extent of abatement required around the existing windows and wall peeling paint. Facilities confirms abatement will be included in the Bid Package for this project. Abatement required is the removal of the interior window glazing putty (with window removal), scraping of the peeling and loose lead paint and removal of all pigeon droppings on the sills and wall surfaces. This will be the same for all four (4) windows in this area affected by the scope of work.
07/27/17 – Controls price came in yesterday for $20,000 to upgrade the system for the new layout. This includes removing three units and extending new controls to the new unit named AHU-53N. ThermalTech proposed new smaller soffit and VAV in mezzanine, county okay with this approach.
08/21/17 – Meeting with TEI and GBBN to view area and existing windows to the south of the visitation booths.
08/30/17 – TEI will get ITB and estimate complete this week, project to be Issued for Bid on 9/7/17.
08/31/17 – BW stated that he thinks the drawings are good to go. TTC stated he needed to update the HVAC details for access through the visitation booth add notes for painting the louvers to match the outside of the building color on that row.
09/19/17 – The prebid meeting was conducted today with good turnout by many contractors, some mechanical and some general contractors attended.
09/25/17 – Addendums 1 & 2 have been issued for this project. The bid opening is scheduled for Oct 12 for this project.
10/24/17 – Addendums 3,4,5 have been issued and the Bid Opening extended to Nov 2 for budget issues and equipment clarifications.
11/14/17 – TEI and GBBN attend Post Bid Meeting with Apparent Low Bidder at TEI’s Offices.
11/28/17 – GBBN and ThermalTech have update the drawings to Permit status and removed all the work not being installed and updated the floor plans to make the fifth floor room the new Mechanical Room. Some work remains and some coordination still remains but we believe it will definitely work it is just small spaces and not a lot of free room around the equipment.
02/08/18 – Const Mtg held. Submittals are coming in and being returned to contractor. ThermalTech met Triton on site for fieldwork.
02/19/18 – All HVAC submittals returned. Permit cost worked out with Triton, they can follow up with the City for costs.
02/20/18 – Door, Frame and Door Hardware submittals stamped “Approved as Noted” and returned to ThermalTech.
03/12/18 – Pre-construction meeting held with Triton for this project. Project schedule, SOV, and additional submittals required from Triton. Permit not ready yet, Triton to work out final cost with city, ThermalTech has worked out all the details with the city plans examiner. Triton to distribute meeting notes and revised schedule.
03/21/18 – VAV submittal returned for Tuttle and Bailey with revised btus per TEI markups. Fire Protection & Controls submittals still need review by TEI.
03/24/18 – HVAC Permit is Issued per EZ-trak. Electrical permit is reviewed and on hold for pickup by EC. Next Meeting is Monday. Triton reported today that the room is not clear for construction yet.
04/23/18 – Const Mtg #4 – Triton is on site working on demo and new work.  AH-53 has been removed from KP area.  New ductwork is being installed.  Existing pipe has been turned off and some of the existing piping has been capped off.  A leak in Cooper area may be related to our work and shutting down our heating hot water to AH-53.  Work is on schedule.
04/09/18 – Const Mtg #3 held and meeting notes distributed. We worked on some of the supply duct routes and return paths.
04/23/18 – Const Mtg #4 held and meeting notes distributed. Reroute return duct KP area, submit CO for sprinkler piping change, finalize HHW and CHW lines to and from units and some of the piping in Shanahan area will be below the duct.
05/07/18 – Const Mtg #5 held and meeting notes distributed. Honeywell on site and starting to work, interior lighting in AHU discussed, const mtgs changed to weekly for final run, Triton to cover all vents in Shanahan’s area and clean debris up daily in work area.
05/14/18 – Const Mtg #6 held and meeting notes distributed. Triton asked to submit change orders for extra work on this project so far. Shanahan soffit will be converted to rigid board ductwork and pants fitting to be fixed before insulation.
05/21/18 – Const MTg #7 held and meeting notes distributed. We are wrapping up Shanahan courtrooms and moving over to Cooper area. Honeywell controls are in place and starting to control the unit. IBI inspection schedule for next week KP ceiling should be going back up, change orders approved, painting has started and we are moving into the final phase of the project.
06/26/18 – Final inspection was approved last Monday. Mechanical and Electrical punchlist have been distributed. Request for Change Order Work #2 has been designed by TEI and sent to Triton for pricing. We have a rough idea of contingency monies left on this project and will try to fashion the changes with that remaining funding.
07/16/18 – Met AAB on site for testing of the airflow through the AHU. We tested the unit at various speeds and took videos on the inner workings of the unit and discovered at 45hz water migrates from the drain pan to the fan section under the drain pan divider. We will contact the manufacturer to rectify this issue.
07/25/18 – Final punchlist was back checked this month. The water leaked returned to which TTE and Triton have been working to resolve with JCI almost daily – there is a lot of behind the scenes work and communications Facilities not included on discussing issues and remedies for this problem. As of yesterday, Triton has increased the depth of the trap, resealed the interior of the unit, and tested interior drainage. JCI tested the levelness of the unit and found it to be installed properly, water was poured into the drain pan and found to drain back and down to the outlet of the unit and at 40hz all seemed to be well.
07/26/18 – The unit was set up to 45Hz and allowed to run overnight at the higher speed. If the unit is still dry we will release it to automatic operation.
08/28/18 – TTC went on site with Triton to work out the reheat coil fitting above the mezzanine in Shanahan’s area.  It’s a very tight fit and the two other VAV boxes have to have their supply ducts move so the new zoning works.  Other than that one coil, everything else was fitting well.  TTE approved controls submittals last week and confirmed with Randy Waddell yesterday for final order.
08/29/18 – Triton reported things are going well but contacted Honeywell to see when they will be on site adding controls and piping to the system.
09/24/18 – ThermalTech met with Triton last week to hash our some piping changes that were not making sense in the field and Triton is making those changes this week, followed by insulator and controls contractor wrapping up.  All work to be done by Thursday September 27.  Triton still working at night.  TEI still looking for hot water reheat lockout temperatures from Honeywell.
10/23/18 – All change order work is complete and the system is programmed into the controls.  Space temperatures look good today and the heating appears to be on.  All spaces are 72 and above.  TEI will start the closeout process on the project.
11/29/18 – Triton dropped off the OM’s yesterday to ThermalTech and TTC hand delivered them to the County including final payapps and closeout paperwork at the Partner Meeting. This project can be archived.
01/22/19 – Project running good. TEI collected all sensors from the space except for the one in the Sheriff surveillance space. Hopefully Barth will recover that for me.
01/23/19 – Project Ready to Archive.

by TEI TEI

COURTHOUSE Roof 10 Repair

Partner Lead: THP

06/11/15 – County requested SOSS.

06/13/15 – THP and County visited the roof.

06/25/24 – THP anticipates issuing the SOSS.

07/17/15 – County Facilities, THP, TTE and Chambers Control met to discuss phasing the project and completing the work.

07/30/15 – Due to the different scope of work, THP split Court House roof project into Roof 10 (repair/replacement) and Roof T (warranty extension) for tracking purposes.

08/01/15 – THP anticipates issuing an SOSS to the County.

08/03/15 – THP issued SOSS to County.

08/07/15 – County approved SOSS.

08/26/15 – Project update: THP developing design documents.

09/24/15 – Project update: THP developing design documents.

10/29/15 – THP completed roof scan, roof cuts and field measurement. Design development documents anticipated to be released 11/2/15.

12/09/15 – THP working on bid documents with update roof configuration from field findings.

01/28/16 – THP working on bid documents

02/25/16 – THP confirmed plumbing routing and delivered update plumbing drawings for review by the County.

03/21/16 – Prebid meeting conducted.

03/28/16 – Addendum no.1 issued.

03/31/16 – Last day for contractors to submit questions.

04/04/16 – Last addendum will be issued.

04/12/16 – Two bids were submitted with both Contractors within the project budget. Imbus Roofing was the apparent low bidder.

04/13/16 – THP completed a cursory reviewed of Imbus’ bid and recommend the County accept all alternates.

04/14/16 – THP issued draft AIA contract for County review.

05/20/16 – County issued PO and Contract to Imbus Roofing.

05/26/16 – Imbus working through background checks and project planning.

07/22/16 – THP contacted Imbus regarding submittals. Ed with Imbus stated he would gather submittals and send to THP.

08/05/16 – Preconstruction meeting completed.

08/25/16 – Imbus in the process of mobilizing.

09/28/16 – Imbus has roof 10 completely removed and the temporary roof installed. Have had some leakage issues, but otherwise turning water. THP provided pipe post support detail to Imbus for pricing. Imbus initially submitted $524 per pipe and revised to $400 per pipe installed. County authorized to proceed with fabrication and installation. Imbus started tear-off and temporary roof installation last night on roof 6. American Façade plans to begin sealant work on cornice adjacent to roof 10 and along north and south walls of law library tomorrow. Imbus and THP meeting tomorrow to review pipe penetration near RTU in southwest corner of roof 10.

01/26/17 – Project update: Roof 10 – complete , Roof 6 – complete except downspout installation, Roof 5 – last phase of layup this week. Flashings, cap flashings, and terminations still to do. Stainless steel trough at Roof 10, painting roof hatch at Roof 5, and HVAC duct replacement at Roof 5 still left as well. Change order for securing electric lines on outside of parapet at Roof 5 being developed. This work is outside our scope, but Imbus subcontractor has provided cost for work. Imbus plans to turn parking in courtyard back to county this week. Hoist and scaffolding was to come down over the weekend. Bert asked THP to look at stained glass windows and protection from winter ice or snow.

02/22/17 – Roof 5, 6, and 10 generally installed. Imbus working on final detailing and flashing conditions

03/23/17 – Project update: Roof membrane installation complete. Imbus has several items to finish prior to reaching substantial completion.

03/23/17 Roof 6 and 10 need punch list developed.

03/23/17 Roof 5 has the following efforts to complete: Reconnect HVAC units, including running new conduit/power on parapet wall. Duct work (elbows) need replaced. Terminations and flashings need to be complete below the duct work. New posts need to be cleaned and painted. New posts need sealant to prevent water from being harbored on top of the cap plate. Prepare and paint the roof hatch. Repair junction boxes. Complete the termination of roof membrane at stained glass window. Repair the CMU block at steel stair landing tie-in at wall. Finish target at roof drain on north side of Roof 5. Fix the walk mats that are not lying flat. Change order work – replace displaced electrical conduits. Finish counterflashing installation.

04/27/17 – Imbus finalizing punch list. Work anticipated to be finished week of May 1st.

05/18/17 – THP issued final punch list.

05/25/17 –  Imbus working on punch list.

07/26/17 – THP to complete review of final punch list.

08/28/17 – Final punch list repairs were reported to be completed Friday 8/25. THP to complete final review 8/31.

09/27/17 – THP to complete final review of final punch list.

10/20/17 – THP publish updated punch list.

10/24/17 – Imbus and County walked the roof to review the punch list.

11/30/17- Imbus completed all punch list items.  County is waiting on final prevailing wage from contractor.

by TEI TEI

COURTHOUSE Cafe Partition Renovation

Partner Lead: GBBN

09/09/14 – William Radcliff of the State of Ohio Opportunities for Ohioans with Disabilities contacts GBBN about a renovation project at the Cafe on the first floor of the Courthouse.

09/16/14 – GBBN completed the required Business Enterprise Price Quote and Form and sent completed forms back to William Radcliff.

09/19/14 – GBBN meets with William Radcliff on site to discuss full scope of remodeling and deliverables.

09/26/14 – GBBN sends final construction documents to William Radcliff as requested.

10/06/14- William Radcliff responds with approval of deliverables.

10/14/14 – GBBN sends Hamilton County Facilities/Purchasing Department Bid Form with language regarding Background Checks.

01/23/15 – GBBN visits City of Cincinnati Permit Center to stamp and sign final permit drawings.

01/28/15 – GBBN sends final invoice to William Radcliff of the State of Ohio Opportunities for Ohioans.

by TEI TEI

COURTHOUSE Asbestos Study

Partner Lead: GBBN

10/20/11 – Kick-off Meeting with Pinnacle and Facilities Department

11/04/11 – Facilities Department sent list of buildings to be included in Asbestos Abatement Reporting.

02/14/12 – Pinnacle Engineering requested drawing files for the purpose of documenting findings.

02/15/12 – GBBN sent drawing files to Pinnacle Engineering in PDF form.

03/30/12 – Meeting with Tim Schue of Facilities Department to discuss Archibus Input Strategy

04/05/12 – Meeting with GBBN and Facilities to discuss drawing format.

04/25/12 – Meeting with Tony Matre, GBBN, Building Manager and Pinnacle to kick-off field investigation

05/10/12 – Meeting with GBBN and Pinnacle to discuss drawing format for reports. Tony Matre’s request for ‘General Information Notes to occur on each floor plan are discussed.

10/27/12 – Pinnacle Engineering sent ACM list from Archibus/BRG for purposes of designation conversion from Pinnacle Report to Archibus entry.

02/26/13 – GBBN inquires with Pinnacle for status of remaining documents to be sent to GBBN.

02/27/13 – Pinnacle reports back that progress has been made and further timetable for receipt will be coming 2/28/13

02/28/13 – Pinnacle sends GBBN a proposed completion schedule for remaining work.

03/18/13 – Tony Matre calls GBBN to request a status check on Pinnacle’s delivery schedule.

03/27/13 – Scheduled to be done April 12.

04/05/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response.

04/20/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

04/25/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response.

05/29/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

06/26/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

07/24/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

08/05/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

08/27/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

02/24/14 – GBBN and Pinnacle Engineering met with Tony Matre to deliver the final reports and electronic copies of the reports. GBBN is completing final report drawings for Pinnacle as final preliminary drawings are delivered by Pinnacle to GBBN. All drawings are to be completed by 2/28/14.

03/26/14 – All asbestos reports and associated drawings are complete. Final drawing input into Archibus/WebCentral is ongoing. GBBN was granted access to the database to complete the drawing downloads. GBBN awaiting final bulk drawing from Pinnacle in order to complete the Archibus drawing data input.

07/01/14 – GBBN sends another email and follow-up voice mail to Pinnacle Engineering inquiring about status of billings and final completion.

08/29/14 – Pinnacle Engineering sends invoices for services to GBBN.

10/29/14 – GBBN sends Tony Matre SOSS invoice status with all of the Pinnacle Engineering invoices included.

10/31/14 – GBBN pursuant to Tony Matre’s request sends letter of documentation regarding scope overruns.

02/20/15 – GBBN in communications with Tony Matre to finalize invoicing for these projects.

by TEI TEI

Courthouse 2010 Service Drive Water Infiltration Investigation

Project updates included with F12.

August 26, 2010 Tony Matre requested option #2 repair (full repair) be completed. To expedite 2010 repair schedule work to be included with F12 as an add alternate. THP to bill design services to remaining hours on F13; then review additional service proposal.

THP delivered report August 25, 2010.

Past Comments

THP delivered report August 26, 2010

THP to complete report.

SOSS approved June 8, 2010.

THP submitted SOSS for review on May 11, 2010.

THP and County completed kick-off meeting.

by TEI TEI

Courthouse – Exterior Lighting Project

04/11/05 – TE still trying to get final paperwork from Mike Cavanaugh. Once we receive this final paperwork, we will send to Mark Hardin.

04/11/05 – ICS completed As-builts and sent to ThermalTech and TE sent to MWH in middle march.

04/15/05 – Received final pay application and closeout documents. Signed and sent to Hamco.

by TEI TEI

Courthouse – Facilities Basement Remodel

Electrical As-Builts are complete. TE is finalizing the drawings for delivery.

HVAC As-Builts are completed. 11×17 color plots have been delivered.

Plumbing As-builts are nearly complete.

Electrical As-Builts are under field verification. We expect to be complete in under 2 weeks.

by TEI TEI

Courthouse – Video Surveillance Project

07/12/05 – Stapleton turned in final pay application, change order and final paperwork. TTC will review this week and get down to the County.

12-20-04 – Construction meeting #2 see meeting minutes.

12-02-04 – Construction meeting #1 see meeting minutes.

11-0-04 – Pre construction meeting took place, reviewed standard county items.

10-05-04 – AIA Contract for Stapleton typed and mailed to Sharon Chamberlin.

Project bid open had taken place. Mike Stiles has contacted Stapleton Rep. and confirmed the package and Sharon has authorized the contract to be prepared. Mike Stiles requested a copy of Stapleton’s bid from JoAnn in purchasing. Waiting for her to let us know when its available for pick up.

01-20-05 – Construction Meeting – see meeting notes below.

03-07-05 – Construction Meeting – see meeting notes below.

04-08-05 – Completed punch list and sent to contractors to complete – see list below.

04-29-05 – Received call from Paul Wilson with CSI. He indicated that the items on the punch list are complete. He also indicated that their time exceeded the bid amount for setting up the NAS device and switch. He request a change order for the additional time. TEI indicated to submit the additional hours and money requested for review by HC and TEI.

by TEI TEI

Courthouse Archive Room Renovation

08/08/05 – TTC received and approved the final paperwork for CH Archive from Ginter. Warranty and other paperwork received. TTC will hand deliver on 085/09.

08/01/05 – Diana called today because only 2 of the closeout documents have been received. TTC is checking with Ginter.

07/11/05 – TE received Ginter’s final paperwork for this job. TTC will review and sent to County.

07/07/05 – TE and PHB worked out the payapp details today by phone. TE had payapp #4 in house but wasn’t logged into the website. Issues are now resolved. Final paperwork will be sent to the county on Friday.

07/05/05 – TE received final pay application for project. There are some details that TE has to discuss with PHB before TTC will submit to County. Owner’s address is incorrect and has been on previous apps – TTC missed that. The previous payments don’t add up. Line 7 (previous payapps) doesn’t match Line 6 on previous payapp. TTC will stop in at PHB in morning and go over with Lee Anne.

06/21/05 – PHB (LeAnn) called to say they are sending a revised final payapp to closeout the project. TTC coordinate final numbers and instructed the to apply for final.

06/15/05 – Dorn Sprinkler Invoice arrived for $220.00 to clean the sprinkler head covered with insulation. TE will mail to Diana today.

06/10/05 – Peck sent a Word document confirming final allowance usages. TE instructed them to sign the change order sent and send back all paperwork. Bernie said he is on it. TTC emailed Dorn Sprinkler about invoice for cleaning the sprinkler head.

05/24/05 – TTC and PHB discussed final paperwork and O&M’s. PHB is finalizing project for closeout. Paperwork should be received in a week or so.

05/11/05 – DES advised to closeout the project and have PHB return the balance of the allowance. TE will write the credit change order for everyone to sign of $1358 and send for everyone to sign. TE also updated the allowance usages for this project below.

04/26/05 – Ginter and Honeywell ran the FA test this morning. Everything passed with flying colors. Larry Keifer asked us to replace one sprinkler head that had some spray foam insulation on it from the ceiling work by JB Schmitt. TTC called Dorn and instructed them to get with Mike Bartlett and take care of the issue ASAP. TTC told Ginter to send in all their final paperwork for the job for approval.

04/22/05 – TTC contacted Bernie about final paperwork and O&M’s for this job and finalizing allowance work too. Bernie said that the would work on it right away.

04/18/05 – Ginter called today to inform us that the city has finalized our final inspection date of April 26 at 6:00 a.m. at the site. JC Schmitt does not believe they need to be there so TE, Ginter, Honeywell and Mike Bartlett will be required to run this test for approval. All the work is already complete and the pretest already passed. Please make arrangements for this test and I believe Jim Ginter already called to let you know about this test.

04/11/05 – TTC called city directly today to inquire about status of permit. It has been two months since we submitted engineering change to this project and TE pointed out that we need get this project finalized. TTC is waiting to hear back from city. City mailed to Jim Ginter on 04/08/05. We need this permit to call for final inspection according to Amit. Jim did you receive the permit in the mail yet?

04/05/05 – Permit Number for Engineering Change to Fire Protection: 2004P07997.

04/0405 – Ginter called and got approval on the permit from Amit Gosh. Once more approval or actually waiting for the city to call us to tell us to pick up the permit – we will be ready to go.

03/24/05 – TTC checked with Ginter today on status. No new information to report – City has permit and is reviewing it.

03/16/05 – TTC checked with Ginter today on status. No new information to report – City has permit and is reviewing it.

03/03/05 – TTC checked with Ginter today on status. No new information to report – City has permit and is reviewing it.

02/22/05 – Ginter reported that all new fire alarm devices are installed in the Archive Room. Joe Dale of Honeywell is working on the new Fire Alarm permit drawings which have to be submitted to the city as an engineering change to the existing permit drawings. Mike Stiles will assist Ginter if needed to submit these plans. Honeywell was supposed to turn in these drawings on Mon or Tues of this week to the city. They are not turned in as of this email and Jim Ginter is calling Honeywell with word that we want them turned in ASAP so that we can pay out the GC in two weeks and have final inspections complete. TE received signed Change Order #002 from Ginter today.

02/16/05 – Ginter reported that new fire alarm devices are being installed today. They will coordinate with building officials on final inspection. MS wrote a change order for Ginter to have Copp provide the power supply, ~$700, that was left off of the door hardware schedule.

02/10/05 – TE has coordinated with the City and Inspector. Ginter has submitted the new fire alarm permit drawings to the city. TE will attempt to keep everyone apprised on the new status.

02/09/05 – GBBN issued punchlist to JB Schmidt for this project.

02/01/05 – Change order #01 for Ginter was issued by ThermalTech to Contractor and County.

01/20/05 – Final Inspection scheduled for February 3, 2005. Building Inspector and Fire Marshal should be present. 6:00 a.m. Larry Keifer requested additional strobes. Te to coordinate with Honeywell to get this work done.

01/18/05 – The Archive room passed its pretest this morning no problem according to Mr. Bob Ginter. Honeywell has to hook up the outdoor air actuator still and that won’t happen until next week according to joe dale (via b ginter). then we can set final inspection. Peck has to modify the outdoor air duct so that it is not connected directly to the unit and blank off penetration at unit as well.

01/11/04 – TE checked on job progress today. Peck reported that unit could not be completely started up because two receiver tanks were missing from the condenser on the roof. Peck found them in their shop and brought them onsite. They will try to have their crew at the building on Wednesday to install the tanks and the associated piping. Peck also noted that the OA duct is connected to the return plenum at an access door that will make unit maintenance difficult from that side of the unit. TTC discussed with Mike and conferred with other TE engineers and there is no reason to bring in 200 cfm of outdoor air for this space. TE will direct Peck to remove the OA connection at the unit and just terminate the outside air duct at the wall with a grille. This will allow a small amount of OA to enter the space to meet the one person 20 cfm requirement by code. If in the future, we need too, we can add an additional exhaust fan to the OA to bring in small quantities directly into the room. This is a better design and use of the equipment in my opinion and the factory startup technician agrees. Please advise TE that this is acceptable to the County.

01/13/05 – Ginter faxed pencil copy of revised Application #1 which has been approved.

01/12/05 – Called Ginter and Spoke with Michael. Michael remembers the problem with App. #1 and didn’t know that the Application was hold and should be resubmitted. Michael thought that it had been sent to Hamilton County and was upset that he would have to resubmit and wait 30-60 days for payment. Michael asked to revise Application #2 combining the Payment Due amounts of App 1 & 2 and resubmit the Application as #1. We’ll skip submission of Application #2 and use the already submitted application #3 which will be on hold for the final documents.

01/11/05 – Ginter Applications #2 & 3 reflects previous payment for Application #1. Application #1 has been on hold since November due to the incorrect change order amount. Contacted Ginter, spoke to Jim Ginter and explained why I could not submit Application #2, he said that I should talk to Michael who is not in the office and that he would have Michael call me.

01/10/05 – Ginter Pay App. #’s 2 & 3 received.

01/03/04 – JB Schmidt is installing the doors and security hardware on Monday and Tuesday and Ginter will hook up the security right afterwards.

12/30/04 – TTC & MB walked the site and MB pointed out that the Danfoss valves being installed may cause more trouble than they are worth as the piping is sold old and the connection to it could result in steam leaks in the piping below the concrete slab. TE will direct PHB on how to proceed

12/24/04 – TTC talked to Peck for and update. The piping and sheetmetal guys should be back on the job on Monday, 12/27. They are currently under the gun at the Quest building downtown and this is pushing our schedule back. TE emphasized that this project is scheduled to be completed at the end of the year and Doug Turner said a significant portion of the remaining work could be completed this week. Hopefully they do make it on site and a lot of the ductwork and piping gets completed before our next construction meeting. TE has to contact Ginter because communications questioned the width of the cable tray Ginter installed. I don’t see making it bigger now, but let me see what they say. All of the remaining HVAC equipment including controls and steam “donfoss” type valves are on site or Peck has them for installation. Mike Bartlett questioned the rubber pads under the feet of the Liebert units. TTC will check into this item with Peck.

11/29/04 – Liebert unit and crane delivery scheduled for this Saturday. TE forwarded request for CH personnel for this weekend work onto the County. TE requested that TE, PHB and Mike Bartlett get together later this week

11/10/04 – TTC coordinated with Peck on the feet of head available from the condensate pump on the Liebert unit and what is called for on the drawings. Pump has 20 ft of head so overhead piping can be ran about 15ft above the floor and along the water line route to the shaft and into the drain per the drawings. AC unit is still schedule to be delivered the 3rd week of November. Peck will still be coordinated the crane and equipment lifts for that weekend. If anything changes TE will let everyone know.

11/09/04 – Ginter Pay Application #1 received with incorrect Chg. Order Amount Jim of Ginter notified and will resubmit.

11/09/04 – Ginter Change Order #1 sent via courier to Ginter for signature and then to Hamilton County.

10/21/04 – Mitch from HCCD at last meeting requested an EMT conduit or cable tray from existing telephone data closet to archive computer room. TEI to get revised sketch to Ginter Electric for pricing.

10/08/04 – TE visited the site. We noted the walls are up to the deck, the caps are on the ceiling cavities per our last meeting. The Liebert HVAC unit is not expected to ship until late in November. Pet.

09/30/04 – Rough in inspection complete for Electrical, Some supports installed, waiting for GC to install doors for door contact work.

09/15/04 – Construction Meeting complete. Project on schedule. See meeting notes below.

09/13/04 – Submittals received from PHB on Mechanical Equipment.

09/13/04 – Permit ready. JC Schmidt to pick up and pay for with reimbursement from PHB

09/03/04 – JB Schmidt on site. Moving studs up into space.

10/08/04 – Sheetmetal caps were installed on existing openings in ceiling.

by TEI TEI

Courthouse – Mediation Department

07/14/05 – Apparently a change order for the VFD drive was never encumbered and this has prevented Reupert from getting paid. Diana said that Mediation was going to work on getting it taken care of.

03/25/05 – Reupert submitted closeout documents and paperwork for final pay application for project. TTC approved and sent to Ms. Supe.

by TEI TEI

CH – Pumped Condensate Project

04/20/05 – All final paperwork arrived for TE approval and submission on this project.

03/16/05 – Process will have final paperwork and payapps in this week.

03/14/05 – TTC contacted Process about final paperwork and payapps.

02/01/05 – Process contacted County to state that project is complete.

01/25/05 – Project punchlist sent out by TE.

01/18/05 – Process has completed pumped condensate work including the summer boiler relief vent. TE to stop by and punchlist on Thursday.

01/14/05 – TE approved 1-1/2″ insulation on all condensate for this project. AKM approved a $1600 invoice for Process to tie an existing summer relief vent into and existing 1-1/2″ pipe with a ball valve. Process is to submit an invoice for this work to Mike Bartlett today for approval so they will not have to pull of the job to do the work.

01/10/05 – Process came on site and began work on new piping for pumped condensate.

by TEI TEI

Courthouse – VFD Upgrade on CT-3 and CT-4

08/08/05 – Final Pay Application was received. Allowance was straightened out. Ginter will keep $1000 allowance and pay DK $680 for additional programming work at the site to set the condenser water temperature to a constant number.

07/18/05 – O&M’s arrived to from Ginter Electric. TE will deliver to County offices on Tuesday by hand delivery.

07/06/05 – TTC delivered final paperwork and O&M’s to the County today. They were hand delivered to Martin.

06/29/05 – Ginter delivered the O&M’s to ThermalTech today. TE will give to the County.

06/16/05 – MB reported that the new cooling tower sequence is still maintaining a very stable condenser water setpoint. All parties are very happy at this time. TE emailed Ginter about final paperwork and punchlist work.

06/06/05 – TTC emailed punchlist to Ginter for final work and sent closeout documents with it for final pay application.

05/20/05 – TTC completed punchlist.

05/10/05 – MB reported today that VFD startup was completed by DK and Stoermer’s Gary John. The County reset the vibration safety switch and cleaned on of the low level sump alarms and CT-3 and CT-4 respectably. Ginter said they would check the wiring if that didn’t fix the alarm issue at the panel. The new program to control CWS/R temperature is in place and working from all first test. As the County brings on all 4 towers

04/22/05 – TTC and Dennis Faul worked on finalizing the cooling tower programming. MB and Mike Denton approached TTC about concerns about cooling tower control algorithms especially where concerned about the screw chiller. To that end, TTC has revised the cooling tower sequences such that a steady state condenser water will be more achievable. It is my hope that this alleviates many of the concerns expressed about this cooling plant. Here is the revised control sequence for approval and comments PDF. This should make commissioning much easier as this control strategy is much easier than the original.

03/15/05 – Programming is partially complete. Testing of program must wait till warmer weather. Project is in a holding pattern until that time. Ginter confirmed today that this was the

case.

02/01/05 – TE ready to punchlist these projects within the week.

01/25/04 – Dennis Faul returned TE’s call yesterday. The DK technician is onboard running the electric for new controls. We cannot test the system until we reach summer conditions but DK is still going to put program in soon.

01/21/05 – TTC contacted Jim Ginter of Ginter Electric in regards to 11,640.00 controls work which was not completed and should not be included in the payment due amount of the pay application. Michael Ginter delivered a revised Application #1.

01/13/05 – Ginter faxed Apps 1 & 2 to TEI for review prior to mailing. Changes required to both. Marked-up and faxed back to Ginter.

01/12/05 – Called Ginter and spoke to Michael. Asked him to resubmit both Apps 1 & 2 with allowance included. Michael to redo & fax copy for approval.

01/11/05 – Called Ginter, Jim Ginter said I would need to talk to Michael who is not in the office and he would have Michael call TEI.

01/10/05 – App 1 @ 22,024.80 and App 2-final @ 1,915.20 rec’d. w/o allowance included in contract price or on Sched of Values. 01/03/05 – Ginter provided a Change /Order deduct of $458 if we allow them to leave existing disconnects in place. They are worried that new dimensions will force changes to existing PVC conduit and this could result in problems with breakage of brittle material. County to advise TE on how we want to proceed.

11/18/04 – TE approved the design of all 4 cooling towers being routed from the main MCC in the penthouse. TE will provide new stamped design drawings for Ginter to submit to IBI showing the new work.

11/05/04 – Pre Construction meeting.

10/27/04 – Contract approved, TE to setup kickoff meeting.

09/24/04 – Mailed Contract to Tony for approval.

Prebid complete. Addendum 1 distributed.

by TEI TEI

CH – Relocate Basement FCU

03/30/07 – DK called today to inquire why they weren’t paid their final monies on the bsmt FCU project. TTC resent the 07/14 data to MWH.

07/14/05 – Received final invoice from Debra Kuempel today. It will be voided because of billing errors.

04/11/05 – TE talked to Michelle Combs today. DK sent an invoice in December for this project. TE does not have a record of this receipt. Michelle will resend a full invoice today for TE approval. TE will send to MWH immediately afterwards.

11/18/04 – Project complete. No punchlist items to complete. All units are running and trades finished up wall and electric work.

11/10/04 – TTC talked to Ed Enderle today. DK is hoping to be 95% complete by the end of the day on Thursday. TTC will visit the site and check on progress. DK is also going to seal up the walls during this project. HamCo will have to reconnect the electric if they haven’t already done so before units are operations

11/08/04 – TE visited the site on Friday afternoon. Fan coils are relocated and piping is in the midst of being changed over. Ductwork has been removed from through wall condition. Wall penetrations are open. Ceilings are partially disassembled for FCU relocation, ceiling tiles are stored inside the rooms.

10/21/04 – Pre Construction meeting complete. Debra-Kuempel scheduled to begin the first week of November. See notes below.

10/14/04 – TTC messed up the pre construction and it has to be rescheduled. Debra-Kuempel and TE will be onsite on October 21 at 8:30 am for the next meeting.

Bidding is complete.

County: ThermalTech is awaiting Notice to Proceed

by TEI TEI

Courthouse/Justice Center – 2003 Boiler Plant Upgrades

11/02/05 – Triton called to say that the certified payroll will be delivered tomorrow. This should wrap up the project completely.

10/18/05 – Triton’s final payapp came in. TTC approved and will give to Tony Matre on Wed via hand delivery.

10/03/05 – The valves and spool piece have been replaced. Triton is getting an insulator down to the site for final work. The final pay application is en route to ThermalTech for final approval.

09/22/05 – Triton is scheduled to replace the defective valves in the Courthouse this week on Saturday. They will being working on Saturday morning. The CH will turn off the steam on Friday to allow the plant to cool down.

09/13/05 – Triton came down and measured for the spool piece today. Mike really wants it installed by 10/01 and Ed Makin said they would try for two weeks. Mike asked about putting back an isolation valve in the spool. TTC said that TE prefers to keep the main wide open with no restrictions plus the pneumatic valve provides isolation.

09/12/05 – Today we kicked of the low pressure steam flow meter project. Process will do all the work during the first week of October. Triton is coming down tomorrow to measure the spool piece for the Courthouse valve replacement.

07/18/05 – TTC contacted Brian Pastura by both email and voicemail today to discuss resolution on the boiler plant valves and back pressure regulator. No return contact as of this email. TTC will continue to follow up with Triton.

07/06/05 – TTC found a quote from JCI to control the bypass valve on the supply to the JC. Based on the Spirax/Triton quote this may be the best solution to the steam supply issue in the

wintertime because the full flow works as proven last year through the bypass. By automating this valve to try to match what the JC needs, we can control total flow and still have the shutdown of the spirax valve to prevent carryover. PDF. TTC sent out Add01 for the Steam Flow meter project. The new bid opening date is July 12 at NOON. PDF

07/05/05 – TTC emailed Triton and Spirax trying to get the repair price out of them for the steam valve and to find out how to independently control the valve without DDC as TTC requested and Mike Bartlett questions (TTC did not have an answer for).

06/15/05 – TTC emailed Alpha Controls about new bid project: Mike, We spoke about 3 or 4 months ago about replacing the existing flowmeter at the Hamilton County Courthouse. I now have to bid out the replacement of the flow meter at the Courthouse. I will need the specs on the valve that you believe we should install. If you remember our final decision was to abandon the valve in the vertical and go with a steam flow valve downstream of the PRV in the low pressure supply to the building. This will allow us access to a large header from the floor with significant straight length. We will provide all the bid documents for three to five mechanical contractors to give us pricing on the job. I will ask them to hire you to provide the meter and do the control hook up. I will need a budget price from you now for your portion and I will need your help on spec’ing the right model number and getting the right flanges and connections in to accomplish what we are trying to do. You will have take the existing control leads and reconnect them to the new flow meter we are providing. As you already know – we don’t need exact lb/hr reading for some sort of process control but rather a flow meter that can give reasonable accurate readings from 0 – 10,000 lbs/hr or so.

06/15/05 – TTC and Pete Rorison talked by phone today. The existing valve will not and can not work under the existing conditions. Spirax can provide a Globe Control valve that can work in this situation. They cannot however, provide this valve at no cost to the County. Pete said the original valve was designed at what they believed was the correct design conditions at the time. Spirax will provide a new control valve, actuator and positioner and re-use the existing pilot positioner tap for control. Either the County or Triton will need to run a new pneumatic line over to the controller/actuator for the motive force to operate the new control valve. TTC instructed Pete to coordinate with Triton and get a best and lowest price together to do this work. There will be a credit for returning the old valve in lieu of the one we may install. When TE receives this new price, we will pass it along with a recommnedation.

06/14/05 – Pete Rorison emailed TE back today. Many of the issues in his memo were not correct. TTC sent a response email to rebut what was said and to try to clear up exactly what Spirax/Sarco is ready to do to make things right. TTC will follow up with Triton tomorrow for final resolution.

06/10/05 – TTC emailed Brian and Pete with a final direction on the Steam Control Valve. TE told them to have a solution in place by June 30th. PDF

06/09/05 – Triton called today to inquire about status of project. Per TTC’s conversation with Tony Matre on 6-7-05 there needs to be some resolution on the Spirax/Sarco valve issue before final payment is received. TTC will call Pete Rorison to see if another valve will work. TE sent out an email to discuss with County PDF.

05/20/05 – TTC and MB discussed the findings of Randy from PHB. The final word is that the steam valve is working correctly. TE next will make a recommendation about increasing the steam pressure in the boiler plant up to compensate for the low flow through the Spirax Valve. Another option is to set the max flow with the bypass open a certain % (field verified) to meet the max demand of the JC in the winter time. PHB stated that the 20:1 spring is the correct spring for the pilot and that the pressures seen on the inlet and outlet of the valve are correct for this valve.

05/10/05 – Triton’s revised invoice arrived. TTC discussed with Jim Drye the steam valve and decided to talk to Rich DeBat, a former Armstrong rep, about this control valve. TTC also contacted PHB about troubleshooting the existing valve. TE had Bernie provide a name of the their steam control valve tech (Randy) who is scheduled to meet Mike Bartlett on Friday at 6:30 a.m. to look at the valve and maybe do some troubleshooting for us. I told him to bill ThermalTech for now (as a reimbursable expense for now) We can judge how much of this time and work we need to help us out. After he gives us a synopsis, TE will provide a permanent fix for the boiler plant and we understand this has to happen this spring/summer.

04/20/05 – Triton and TE discussed in great deal when Triton is going to get paid for their remaining outstanding balance on this project. TE sent an email to Tony about what Triton wants to do – bill the whole contract now at this point. We are waiting on the County to decide about this.

04/19/05 – AKM & TTC talked about how to resolve PRV issue and Tony said that he has $10,000 to get these final two issues resolved through Triton Services. They will charge to the existing PO’s that Tony has available for this job. TE promised to have these issues taken care of this summer before winter ever gets here.

03/25/05 – TTC and County discussed solutions for the PRV valve since Spirax cannot provide a valve big enough to handle our load. TTC wants to bring in another engineer to discuss placing a second control valve in line to meet the maximum demand and still protect the Courthouse from steam spikes of the Justice Center.

03/14/05 – Triton delivered all final paperwork to ThermalTech for this project. TE to review and deliver to County.

03/11/05 – ThermalTech, County, Triton & Spirax/Sarco met to discuss the regulator valve issues at the CH. Several problems were discussed and solutions were also discussed. The bottom line is that the valve is not sized large enough to handle the full load required at the JC and still maintain the steam pressure required at the JC. Per Mike Bartlett’s excellent suggestions, TTC is pursuing JCI about automating the bypass valve to make up for this discrepancy. More info is coming in the near future to further discuss this issue.

03/08/05 – TTC has spoken to Brian Pastura about getting chuck on site to make these changes and they will do what they can but would like to have a meeting between TEI, Mike, Tony and Triton to decide on how to handle the workings of this valve for now and in the future. i agreed with brian that this would be a good idea and asked him to set up this meeting.

03/07/05 – Mike reported that Alpha’s service called yielded some design and equipment problems with the existing flow meter: Mike Maloney from Alpha Controls came in Thursday to check the Courthouse Yokogawa Steam Flow Meter. He found the sensor not working and recommended replacement. He estimated a replacement cost of just the sensor at $1500.00. After looking at the meters specifications, I found that this meter does not read steam flow under 3000 lbs/hr. An 8 inch valve is set up for reading between 3000 lbs/hr to 40000 lbs/hr. We believe that the Courthouse steam flow would be no more than 10000 lbs/hr. He recommended downsizing the valve to 4 inches and redoing the piping to allow for more straight piping that is recommended. He says that right now the 8 inch meter does not have enough straight piping for correct flow readings. He estimated a complete 4 inch meter at $2000.00. TE recommends doing this work to get this additional flow added to the Master flow readings on the front end for use and troubleshooting reasons.

02/24/05 – Pete Rorison called today to basically say that ThermalTech gave him the wrong design conditions for the valve and that it will meet the design conditions of 65 psig inlet and 50 psig outlet at 16,000 pph. The problem is that the JC suffers and 8 lb pressure drop at full flow so the valve should be sized at 65 psig inlet and 58 psig outlet. ThermalTech sent their rebuttal and the Submittals data stating the the valve was to operate between 50 – 80 psig.

02/15/05 – TTC contacted Triton and Spirax about pressure regulating valve and its proper workings. We will advise when we have more information.

02/01/05 – TE contacting the manufacturer about valve failures. Plants scheduled to be shutdown and worked on when OAT is above 50 degree F.

01/25/05 – Triton tried to fix the pilot positioners but the Powell Steam Valves did not hold. TE requested that Triton contact the mfgr and get new valves to replace the defective valves. The spec and the contractor wanted positive shutoff valves in this system. MB may talk to JC about a steam plant shutdown to replace the valves and do the pilot control work on the steam pressure regulator/backpressure control valve.

01/14/05 – Tim Cancilla is supposed to be onsite today to readjust the controller valve for correct steam flow and to rectify pilot tubing installations.

01/12/05 – Called Tim @ Triton to verify final doc list rec’d. Tim said finals will take awhile to get together and that he needs a Affidavit of Contractor Prevailing Wage form (2 Affidavits mailed), App 5 sent via courier to AKM.

01/11/05 – Revised Triton Apps 5 & 6 rec’d, 6 on hold for final docs, 5 signed.

01/10/05 – TTC request I notify Triton of incorrect contract amount on Apps. Tim w/Triton notified that Apps 5 & 6 should be resubmitted w/correct contract amount & that 6 will be held for final docs, refaxed final doc list.

01/04/05 – Tim with Triton, faxed a pencil copy (Application No. 6) requesting $27,206.76 with zero balance. Pencil copy approved.

01/04/05 – Email AKM regarding status of Triton App 5 pymt.

01/04/05 – 010405 App 7 received w/zero balance ( Requested 9,571.00 Change Order amount). App 7 voided due to incorrect $’s on G702 and G703, Called & faxed mark-up of 7, requested a revised App be sent as App 6.

12/10/04 – Follow-up with Triton Services, Timothy Herzog regarding final documents. Mr. Herzog requested that ThermalTech re-fax the list of required documents (which was done in the am of this date).

120605. T. Herzog contacted & fax sent informing that App 6 on hold for final docs.

11/22/04 – Triton Services, Timothy Herzog contacted regarding final documents and a list of documents required faxed to Triton.

11/16/04 – Process has replaced the traps and they are insulating them. Kerry approved contract changes and TE will deliver the documents to Tony for final work, Triton will bring diamond plate down for installation and have their insulator finish up the piping on the boiler room. Mark Neary returned my phone call and I have to follow up with him on flow meter recalibration.

11/04/04 – TE contacted Triton about finalizing the punchlist on this project including the diamond grate, insulation, flow meter, summer boiler vent line. Process has been told to replace the steam traps asap even pending the outcome of the purchase order submitted. TE instructed Doug Doxsey that in addition to the traps and the pumped condensate lines that a steam trap to drain off the summer boiler may also be a change order the County is looking to get prices on.

10/27/04 – TE sent change order request to Process for steam trap replacement. PDF

10/22/04 – TE scheduled to perform punchlist on project. Insulator has wrapped up majority of work. Steel grating not installed yet.

10/21/04 – TE has to get a steam trap decision to Mike Bartlett this week. TTC has to send data to Michigan to get answer from our special Armstrong Engineering.

10/21/04 – TE and MB discussed having process add the summer boiler vent to the new venting scheme instead of floor connected.

10/20/04 – Three Original AIA 101 Process Construction Contract document and Original plus two copies of submitted contract documents, from Process, sent via UPS to Hamilton County Dept. of Facilities.

10/19/04 – Original Contract documents received from Process Construction.

10/14/04 – TE approved Change Order #2 and sent 4 copies to Brian Pastura for Triton signing and then onto Tony Matre. Also the insulator is making steady progress on finishing the rest of the boiler plant insulation work. Once complete TE will do final punchlist.

09/29/04 – Triton is starting on concrete work, Neary provided controls programming earlier this week and appear to have done a nice job, Mike Bartlett reported that a recent Justice Center steam spike of 7000# reduced JC pressure from 55 to 14psig but the CH plant went from 55 to 51psig thus the new regulating valve appears to have worked as the plant maintained nominal working conditions. TE to schedule their punchlist as soon as insulators, concrete and controls are complete by Triton.

09/24/04 – Tony approved bid for Process at $9680 for pumped condensate work. TE will write contract.

09/24/04 – Water hammer is occurring at the CR-2. Stoermer-Anderson and TE will have to study this scenario and advise the County as to what might be occurring and how to eliminate this effect. Theoretically with no “live steam” going to the receiver, we should not have steam hammer at this piece of equipment.

09/24/04 – Triton submitted the NO BID paperwork so the bid tab is correct. How should we proceed. Please advise.

09/23/04 – Triton updated TE on project status today. Concrete contractor should be done in the next couple of days. PCI should be onsite today starting insulation. Ed Makin will be back down to put the bypass around the city water to the CR-2 on Friday maybe.

09/23/04 – Bids are in for pumped condensate work. Process is low with $9680 bid. Bid Tab – PDF

09/13/04 – Triton has completed work inside the boiler room and removed all tools. Room has been cleaned in general. Outstanding items include: Controls work, final inspection, concrete work, lead/lag switch on CR-2, plumbers must fix softeners so that we can start boiler plant, air control regulator must be replaced, insulators must do their work, steel grating must be installed at entrance to CR-2 room.

09/08/04 – Inspectors questions flash tank drain going directly to floor drain. TE to coordinate in morning on 10th. Code states there must be an air gap of 1″. Existing condition did not have it installed that way. TE to resolve.

09/03/04 – SA instructed us to take the new 3″ pumped condensate into all 3 of the water inlets in the top of the tank. This is a changes from the design documents. Triton will have to submit a change order for this work.

09/02/04 – MB reported today that Triton installed the 3″ make-up from CR-2 to DA into a 3-inch opening without a spray head. After much discussion, it was decided that Triton has to install the piping such that the 3″ CR-2 feed must tie into (2) existing city water connections both using spray heads. The remaining opening can stay in place for an emergency city water feed control to the DA.

09/02/04 – TTC and MB decided that after the plant is complete, we will begin to run the plant at 65 psig just like it has always been designed to operator and therefore we won’t need the PRV’s installed at this time.

09/02/04 – Dorn contacted TE today. TE gave them a PO that will be reimbursed during subsequent invoice from TE to do work on moving sprinkler line on T&M basis in CH walk-in cooler.

09/02/04 – Mike Bartlett report that Spirax stated that strainers must go in front of the PRV valves. I checked with Todd and Gary and the HTX are rated for 125 lb service and they should be able to take the higher pressure in the boiler plant. I am recommending that instead of incurring more cost, we just return the valves and take the credit to the job to offset some other change order cost on blowdown lines.

09/01/04 – TE talked to Spirax engineers. Both PRV’s must be installed for HTX’s, flow meter can go in vertical, they will provide flow differentials based on varying steam pressure.

09/01/04 – Stoermer Anderson confirmed that summer boiler relief valve is already set at either 80 psig or 100 psig, so no additional work is required.

Triton Services will be back on site August 16 to begin finalizing the project work

CR-2 and DA piping will require a FULL plant shutdown scheduled for August 20th.

Summer boiler must be online independent of the main plant by August 20.

TE will coordinate with Stoermer Anderson for final Summer boiler startup details

Summer boiler will use treated city water during 3 week shutdown as work is completed.

TE and Triton to coordinate schedule with Mark Neary for new control sequences and SA for boiler work and equipment changeout and use.

County will provide an airline close to top of CR-2 so that project doesn’t incur additional change orders

Boiler plant work scheduled to be finished and back online by September 10, 2004.

Vendor: Triton Services | Date Issued: 08/23/04 | Date Answered: 08/23/04 Question: There are (2) 3″ inlet to the DA tank, can we remove the 2″x3″ bushings and run the pumped condensate directly into the DA tank. Response: The bushing and utilize one 3″ inlet for the 3″ pumped condensate from CR-2. Since all material are now steel, the existing dielectric flange for the brass fitting will not be required. Keep one of the existing connections in place and utilize the existing city water connection and controller as backup or emergency fill if ever needed.

Vendor: Triton Services | Date Issued: 08/23/04 | Date Answered: 08/23/04 Question: Triton asked TE to locate the circuit setters shown on the drawings. Response: TE advised Triton to NOT install the circuit setters shown on the drawings and save the money to use on the blowdown piping.

Vendor: Triton Services | Date Issued: 08/23/04 | Date Answered: 08/23/04 Question: Can 6-inch piping to steam regulator on catwalk be maintained in size. Response: TE and Triton met onsite to discuss field issues and many small items were resolved including the piping layout for the new back pressure regulator valve on the catwalk service the JC. 6 inch piping can feed the JC and increase to 8-inch immediately after the final isolation valve just like existing piping is installed. Existing Justice Center main will now be bypass. Per the contract documents Triton will replace regulator on wall with new equipment to operate bypass valve.

Vendor: Triton Services | Date Issued: 08/23/04 | Date Answered: 08/23/04 Question: How do the new control valves get place and installed on the project? Response: The new control valves for the CR-2 and DA-1 shall be installed per the direction of Stoermer Anderson. The existing city water feed to the DA can remain in place as on 3″ outlet can be used for the feed from the CR-2 unit and the other 3″ (bushing down to 2″) can remain as emergency feed to DA.

Vendor: Triton Services | Date Issued: 08/24/04 | Date Answered: 08/24/04 Question: The original separator location was in the Justice Center. Now the separator is being installed in the Courthouse. This is a change in the contracts. Please advise. Response: Brain, you are correct the original location was in the Justice Center but it was to be installed over 15ft high in the air requiring lifts and extra manpower to reach it and install it at the location shown on the drawings. All of this work is included in your base bid. It is my belief that the new location provides a significant reduction in manpower and installation time. Therefore I see no reason to incur additional cost to the project for this project change as we can just use that savings to apply it to the additional work to reroute the condensate line which should only include 2 90 degree elbows and their associated welds. The rest of the scope remains the same.

Vendor: Triton Services | Date Issued: 08/25/04 | Date Answered: 08/25/04 Question: There is one 3″ flange on top the DA that has significant pitting. Should this be replaced? Response: Please replace the flange at the weld neck. Make it suitable for the blind flange required in the construction documents.

Vendor: Triton Services | Date Issued: 08/25/04 | Date Answered: 08/25/04 Question: The separator has a 1-1/2″ outlet on it. The TE drawings show a 1″ outlet with a 3/4″ condensate drain. How should this be remedied? Response: Please route the condensate drain as 1″ to the condensate system.

Vendor: Triton Services | Date Issued: 08/26/04 | Date Answered: 08/26/04 Question: What pressures should the new flow meters be calibrated too? Response: Set the Courthouse flow meters to 65 psig and the Justice Center flow meters to 55 psig. Name them CH-FM3 and CH-FM4 and JC-FM1 and JC-FM2

by TEI TEI

Courthouse/Justice Center Separator Installations

05/04/06 – Tony called today because Geiler’s final payapps arrived unsigned by ThermalTech. TTC drove down picked them up, signed them, made copies and sent them back down to the County for payment. Geiler did not sign the copies so TTC just copied the original after he signed it. Once paid this project should be complete.

04/10/05 – Indrolect inspection at the Justice Center is finally complete and approved. We are now ready to finalize and archive this project.

02/17/06 – I don’t think the inspection will be next tuesday because indrolect just got to the bottom of Geiler’s permit applications. apparently geiler took out two permits even though they didn’t know it. the city doesn’t even have a copy of the jc one in their file but on the computer it says it was approved. ibi will accept the previous plans exam as valid for the work but now has to schedule another field trip to sign off on the jc wiring install. when ibi decides they can come out, indrolect will call us and let us know and I will call you and let you know.

02/114/06 – Dave Memory of IBI approved the CH install with TTC, MB and Don Ritter in attendance.

01/11/06 – Geiler and Indrolect have submitted for all permits. As soon as plans exam is finished, Indrolect will call for final inspection. Should be within a week or so.

01/01/06 – The job cannot be closed out because of permitting issues from the contractor. Indrolect and Geiler have been given the preconstruction meeting minutes where we state that permits are required on this project. TTC sent Geiler an email today saying that the additional $1024 in allowance money that is being paid to Geiler should be used to acquire all the permits on this job. TE will provide additional stamped sets of drawings as necessary for permits.

11/28/05 – TTC received all final paperwork and will deliver to Tony Matre on 11/30/05. This job is complete.

11/03/05 – Lisa Huber said that all paperwork has been sent to ThermalTech for this project. Certified Payroll will go directly to Tony after they get it on Friday.

11/01/05 – Finalized final billing strategy for final bill. Tony said they don’t have any certified payroll and they need final pay applications. TTC emailed everyone to let them know.

10/20/05 – Lisa from Geiler returned TTC’s phone calls about project closeout. TTC returned on 10/24 at 5:30 p.m.

10/18/05 – TTC emailed and faxed the change to Michelle Wells again in the hopes the final payapp will soon be processed.

10/14/05 – TTC contacted Geiler again for the correct final pay app. No work yet.

10/03/05 – Still no word back from Geiler on corrected pay application.

09/21/05 – We received the final pay application for this project as application #2 this time. This payapp is not right because of the change order amount is wrong versus the allowance. TTC contacted Geiler to advise them on how to fix it.

08/08/05 – TTC reviewed and will deliver the final paperwork to Tony on 08/09. We have not received a final pay request nor the signed punchlist. TTC emailed Barry Bowman to inquire about final pay application.

07/25/05 – TTC received in his mailbox the final paperwork.

07/18/05 – Contacted Barry Bowman about finalizing this project and sending all paperwork in.

06/10/05 – Relisted Punchlist PDF.

05/20/05 – After cleaning the system several times and removing large amounts of rust from the piping inlet to the separator and disassembling and reassembling the 4″ supply line to the separator, the team was finally able to start the separator and see significant flow at the cooling towers. The system when new provide 40 psig of gauge pressure at the outlet. TTC stated that with this much pressure and flow out of the nozzles that 2 towers could be scrubbed at once and the normal operation would be to run the separator with CT-1, 2 for a week then switch it to CT-3,4 for the next week. The punchlist will be sent but Geiler took care of issues at the site during startup. The strainer’s and filters should be changed a couple of times a day for the week or so. Monitor how much cleaning the SEP-1 unit provides and modify the cleaning schedule as needed.

05/17/05 – Startup for CH separator is scheduled for 05/20/05 at 8:30 a.m. with TE, Geiler and Trane.

05/16/05 – MB emailed today to say that all four CH cooling towers are online and ready for separator startup. TTC emailed the contractors and requested the dates. TTC approved application #01 for payment and us mailed to Tony Matre.

05/02/05 – Geiler promised that Bob or his associate would be at the JC today working on the punchlist items for the JC separator.

04/22/05 – TE contacted Geiler for a third time about completing punchlist work at JC. HS confirmed again that no punchlist work has been done. Geiler wants paid badly and keeps inquiring about payment and TE keeps pointing out deficiencies in payapp procedure. TTC believes that we have it resolved now. CH staff believes that their towers will be online next week for startup and testing.

02/11/05 – Punchlist distributed to Barry, Tony, Howard and Mike.

02/10/05 – Geiler asked to see the original meeting minutes where we told them to put a drain and valve into the Courthouse Separator system so it could be used in wintertime operation. TE to send out from original pre construction meeting.

02/01/05 – TE ready to punchlist these projects within one week. Punchlist is complete. TTC to distribute.

01/25/05 – Indrolect made the tie-in into the bus bar on the elevator equipment room yesterday without incident. We powered up the disconnect and then removed the fuses awaiting startup by Geiler. TE contacted Barry about status of project and if it was ready for punchlisting.

0/1/21/05 – No response to the previous day conversation. Followed-up with a fax to Geiler about Application #1.

01/20/05 – Pay Application #1 placed on hold. The Application was not signed or notarized and also did not include a completed, notarized, Affidavit of Original or Subcontractor. Michelle Wells contacted and told of the reason the pay application would be put on hold. She asked for a voice message to be left for Lisa explaining why the application is being put on hold (which was done).

12/03/04 – County approved Geiler change order allowance usage for redesign of PVC system for separator. $3976 will come out of the allowance on this job.

11/22/04 – TTC stopped at the Justice Center this morning and reviewed the existing piping layout with Joe Phillips, Howard Seal, Barry Bowman. What we are proposing to modify the design so that each tower can be isolated from the system and cleaned to prevent any of the “dirt” from the system from getting entrained into the condenser water supply to the chillers. Each separator could be piped separate with PVC and ran an determinant amount of time to clean the tower daily and then return to normal operation. TTC will provide a revised drawing to the Geiler company for this work today.

11/12/04 – TE contacted Geiler about completing the Courthouse and redesigning the Justice Center to meet actual design request. TE will follow up with Geiler.

11/04/04 – TE contacted Barry about field progress on this job.

10/20/04 – Geiler needs to install a drain and valve on CT-2,3,4 mains for winter draining. Install a drain on the main unit. Possibly relocate the existing control panel 90 degrees rotated to meet NEC requirements.

09/29/04 – CH CT 1 & 2 piping is complete. CT -3 & 4 and JC work to be scheduled in the upcoming weeks.

09/24/04 – Geiler estimates that the courthouse should be done is 1 – 1.1/2 weeks.

09/23/04 – TE emailed Geiler to inquire about project schedule update since progress is slow.

09/13/04 – Piping is being installed in Courthouse cooling tower #1. Looks good.

08/01/04 – Geiler asked how do they fit piping under fill? on 08/10/05, TE advised to decrease manifold size to 2″ and slide piping under existing fill. This will maintain integrity of cooling tower. TE provided a new highly detailed dwg of the tower piping for Geiler to use on all towers for this project.

Geiler has been onsite doing the work.

ThermalTech received written approval for sweeper design.

2″ pipes will be slid under existing fill.

Contract should be back on site soon.

Cliff Wingerter is on medical leave so all questions should be routed to Barry Bowmen.

by TEI TEI

Courthouse – LIP Schematic

10/14/05 – TTC has inquired about status of project. If the new door is being handled within, then this project is complete. Diana and Mark please advise as to the status of this project.

05/12/05 – We all discussed alternate schemes for a new door to meet state and local requirements and TTC sent this color scheme with 2 options to GBBN for review and for GBBN to forward to the County for approval. The County shall make the final decision as to which way they want to proceed. PDF

04/26/05 – TE recommended provide a new door in the space to meet code requirements from state and city per this PDF. TE must wait for County to approve a way to go on this issue before project is complete.

02/2/05 – TTC talked to Mike Bartlett and found that complaints have subsided since Greg greatly reduced the total amount of air being fed into the space. This makes sense since the space was being overcooled because the thermostat was on the second floor in the courtroom. Mike and I agree that this should work until warmer weather where the airflow into the space will have to be increased and further substantiates the need for a revised VAV system in this space. As it right now, Mike and Greg are acting as a manual VAV system by increasing and decreasing air into the space as needed. This also shows us that if you are really crunched on money, you may be able to only put a couple of VAV boxes into the first floor area and begin to achieve some control with the hopes that a full redesign can be done later.

02/25/05 – Client reported that space is still not conditioned during the evening time properly. The daytime loads can still never be fixed without a new VAV system and HW reheat (see SOSS TE-115)

02/10/05 – TE has not seen a balance report or payapp for the additional work.

02/01/05 – MB reported today that Room 169 reported no comfort problems with the space except when the main ahu turned on. PHB still has to balance the new air device. TE to punchlist this job shortly.

01/31/05 – MB reported that PHB finished the job last week. TE has to confirm. We will not be able to assess whether this project was a success as of yet because the units are being operated 24/7 in that area due to other heating and cooling considerations.

01/24/05 – Peck started the job. They need a different transition and will be fabricating that in their shop and finishing the job in one or two days.

01/18/05 – Peck said they should start on Wed or Thur, Jan 19, 20 per Ken Shad.

01/03/05 – Peck should be onsite at the end of the week or beginning of next week to start after hours work.

12/24/04 – TE has given Peck a scheme to price up. Assuming they have the best price, they are the odds on favorite to do the work since they have an unused $5000 allowance at the CH in the Archive Room.

12/10/04 – Price came back from B&J at $6000, TE is now inviting Peck to give us a price since they have allowance money left on 2 County jobs. TTC will bring down Peck for a price quote next week. We will coordinate the visit with Mike Bartlett.

11/05/04 – TE met with Mike Bartlett, Bill Scholl & Jim (Bailiff’s) and reviewed the general layout of the 5-ton split system into the Bailiff’s area but the actual unit would be in the Room #167 space (this would have to be approved by LIP), ductwork would have to routed above the 167 space and over to the 169 area, 6 diffusers at 250 cfm would have to be installed to meet the 2000 cfm requirement of the 5-ton unit with a return ductwork. the final design is still under review. TE has asked B&J Jacobs to take a look at the design to provide their opinion about what is easiest and most cost effective. TTC will provide prices associated with each of these schemes so that the County can make an informed decision.

11/04/04 – TE discussed a ductless split system installation with Mike Bartlett, Bill Scholl, Diana Supe and left a message for Tony Matre in the 24 hour bailiff’s area today. TE will follow up with Tony about funding this project ASAP.

10/25/04 – Bill Scholl called today. A full size duct won’t fit through above the ceiling into the bailiff’s area due to a water pipe on the other side of the wall but a duct sized at 12″ x something will. I told them to find out and maximize the size of this duct and proceed with the design. We are trying to get as much air into the space as possible without spending much money. This is the first stage of the “fix”. Other things may have to be done later to enhance the space control.

10/14/04 – Mike Bartlett showed TTC B-12 which is actually the unit that serves the Concessions area. This unit is served by both heating hot water and chilled water. There is a temperature sensor in the supply duct outlet that monitors the leaving supply air temperature. TE believes that this is the sensor that Bill Scholl was talking about in our 10-12-04 meeting. When there is no building chilled water, the DX unit provides cooling for B-12. Our drawing and scheme will still work as the timeclock will divert air to the room #169 space starting at 7:00 p.m. – 6:00 a.m. and regardless of how B-12 is providing cool air it will get into 169 to assist in cooling the space. The system should still be modified as design on the ThermalTech packet.

10/12/04 – TTC met with DES, MB and Bill Scholl to discuss the reworking of the ductwork on unit B-9 to included Room #169 during the evening time. Trades and HVAC will do the work in the next couple of weeks. The plan will have B-9 provide cooling to the space after hours between 7pm and 6am. TE also stated that the ceiling work required should be done after hours as this will be a messy job especially in the kitchen area above the sink. We will monitor the temperature during the first few months and see if the County needs to add additional controls to the new air damper later. Bill reported that a Paragon timeclock will be used as they have one of those instock and the damper and actuator have been ordered. TE is fine with this information.

10/08/04 – TEI received a letter asking a question from the plan examiner. MS needs to talk with him because a letter will answer his question. However, he is out of the country till the 18th. They directed me to his supervisor Bill Spurling. TE have been playing phone tag with him for the past couple of days.

10/07/04 – TE visited the site and TTC has sketched up a possible solution to the Room #169 heating issue. TTC will deliver to the County and look at making the changes utilizing existing equipment for the work on 10-08-04.

09/13/04 – Fire Test at 6:00 a.m. Space did not pass. Egress into LIP has been questions by Larry Keifer. Security drawing was not included in permit set to City. Mike Stiles to rectify ASAP.

09/03/04 – TE delivered a set of permit prints to Brian Meyer for use on this project.

Security issues are ongoing between LIP and Recorders.

Trades doing all work on this projects.

10/18/04 – TEI received question from plan examiner and TEI responded. Plan examiner was out of the country until Monday October 18th. TE waiting for response.

4/19/05 – TEI recieved call from Mark Harden indicating that the State Agency that governs the computer system information will not accept the security level provided by the rear door in the computer room. However, the motion sensor that was added was per the the state building code. Mark and TEI to discuss options. TEI contacted Ben with GBBN to discuss the option of another 2nd path of egress. Ben to respond.

by TEI TEI

Courthouse – EMU Relocation Renovation

10/14/05 – Wayne Trent advised that some portion of this project is moving forward, TTC and Spencer are trying to find out what part of scope is confirmed.

10/03/05 – Spencer is still working with County to get project scope finalized.

08/23/05 – TTC and MS met Spencer Johnson and Mike Bartlett at the site to get field work to prepare for the permit drawing due at the end of the week.

by TEI TEI

COURTHOUSE – BREAKROOM REMODEL

12/26/06 – The project is complete and the work was completed by EDI. The only problem is getting the city buy into the final design. A problem exists between Steve Ossege and ThermalTech about what engineering dwgs were submitted on the project. TEI gave a engineer dwg to EDI to put the work in but the city permit doesn’t reflect it. Once this resolved the project can be closed out and archived.

12/04/06 – EDI has installed the drain pan under the FCU, TTC needs to resolve the dwg issue with the city examination department and get Steve Ossege approval on the closeout.

11/20/06 – TEI has been working with EDI to get Steve Ossege to approve the building permit for this space and rectify the deficiencies in this space according to Steve Ossege phone call to TTC at TEI about the space and the permit. The ductwork has been modified so the return is fed back tot he unit, Mark Donnelly’s office has an OA duct and a supply duct and the FCU will get a pan underneath it as soon as the water switch arrives.

05/26/06 – Dorn received building permit number from Spencer today and is applying for actual fire sprinkler permit work.

05/23/06 – TTC contacted Dorn sprinkler about the progress of the fire permit for the breakroom area. Dorn sent a dwg and a revised dwg to TE showing the changes. Bert Watts will inform us when the ceiling grid is in place and then Dorn will put the new sprinkler heads in the new space. This should work fine.

03/06/06 – TTC got a late start on the project that was supposed to be done last Wednesday but today finished up both the mechanical and plumbing but realized a big problem with the HVAC and asked for guidance from HamCo. TTC sent an preliminary HVAC plot to the County for review. The plumbing is finished and ready for permit.

by TEI TEI

Courthouse Electrical Device Identification

01/10/07 – Completed in 2006

1/08/07 – NR picked up the panel index binder from Hamilton County Facilities. There are no open issues left on this project. All documents have been turned over to the appropriate parties.

1/03/07 – NR met with Jim Eby of Hamilton County Facilities to review the as-built drawings for this project. The open items were resolved. (3) full sized sets of drawings have been delivered to Hamilton County. The binder containing each panel index is in the possession of Bert Watts. A copy of this binder shall be sent by Bert to ThermalTech Engineering.

1/9/06 – 2nd Floor through Roof is complete. Presently working on the Basement & 1st Floor

12/15/05 – 4th Floor through Roof is complete. Presently working on the 2nd & 3rd Floor

11/15/05 – Roof is complete. Presently working on the 5th Floor

by TEI TEI

CH – 2nd & 3rd Floor VAV Renovation Project

03/15/07 – TTC has received and printed out as-built drawings. The entire packet will be delivered to the Courthouse on Friday.

02/05/07 – EDI sent a revised payapp #7 in that looks fine but they also are asking for additional monies for ceiling work.

02/02/07 – EDI’s final O&M’s checked over and are fine. EDI stated that As Builts are included with them. TTC to confirm. TTC rejected pay ap 08 because it shows a negative balance due on this project.

01/22/07 – All Closeout documents received. TE set to work on as-built dwgs to turn over to County. O&M’s and closeout paperwork and pay applications to be sent directly to Tony Matre on 01/23/07. Project was archived by TTC on this day.

08/25/06 – O&M’s have been received. TTC will drop them off to Tony today by hand at the Elevator bid opening. TTC will schedule a punchlist for the job next week

08/15/06 – Payapp # 5 was received by TEI.

07/25/06 – Mike emailed today to report that balancing is supposedly complete. Honeywell needs to finish up because tenants are complaining about space conditions. TTC forwarded to EDI and will get the project completed.

05/22/06 – TTC talked to Mike and John both of whom like the proposed changes to the HVAC system for the courtrooms on the second floor of the new space. TTC will advise the EDI team that we can move forward with the revised layout. Here is a copy of the revised changes PDF

05/16/06 – TTC contacted EDI about change order on pay apps and about next const mtg with john and mike at the courthouse.

05/02/06 – TEI set a construction meeting between TTC, MB and John from EDI for tomorrow to go over project status.

04/11/06 – Const Mtg is tomorrow, TTC to have new change order drawing ready for EDI to go over changes in the space per TTC’s discussions with Mark Donnelly today at the CH site.

04/06/06 – TTC typed and sent out Const Mtg #2 notes. Mike called about fire damper locations. TTC responded that none should be required but if EDI found a certain fire wall we are penetrating that TEI will answer those conditions on a case by case basis.

03/16/05 – TTC wrote change order for EDI to do FCU relocation in basement.

3/15/06 – TTC talked to the Controls Contractor today to review the submittals data. Honeywell will field locate power supplies and provide quantities required for complete installation. They will coordinate with Mike Bartlett for power supply locations and update EDI as-built drawings for future record. Final questions was how many VAV boxes. We confirmed there are 17 new boxes. 9 on one floor and 8 on the other. The Computer room VAV is existing.

03/09/06 – Const Mtg #2 was held yesterday and the meeting notes distributed today. EDI emailed the employee list to Tony today.

02/27/06 – Pre construction minutes distributed. Submittals have been returned.

02/16/06 – TTC approved using NEBB approved balance reports on this project and in particular the sub-contractor Approved Products. TTC will send out the meeting notes later today to the contractors and owner.

12/14/05 – Contracts came back, TE edited, corrected, printed and sent back down to County for commissioner approval.

10/24/05 – Final Addendum went out on job. Other control contractors were added by Addendum. TTC answered MB’s questions about reheat temperatures in the coil locations in saying that coming off the end of the HHW loop is okay, we just need to have the coils selected properly at the time of purchase. 10/14/05 – Sam Boehringer will be handling the prebid mtg for this job as TTC has a traffic court date with his wife at the same time and day. TTC will join the party after court. Habegger called and will be added to the acceptable manufacturer’s list for this job.

10/13/05 – The project has bee publicly advertised. The ITB is 160-05. Bid Opening scheduled for 10/28/05. This is a strictly mechanical only project with Honeywell or Habegger as the controls contractor.

06/16/05 – TTC & SBO met with Diana and Mike to discuss possible scopes for this project. TE designed the full system with all VAV boxes and reheat boxes. The new cost is significantly higher with a full zoning and reheating plan ($161,000). An alternate plan that may work (only 8 boxes) was also given to Facilities ($51,000). It was decided that to half do the design would only waste money and inconvenience the tenants twice. Diana will approach Tony about the budget and doing the full job.

04/04/05 – TE and CH Facilities had first design meeting discussing scope of work and project goals. PDF

by TEI TEI

Courthouse – Blowdown Project 2006

07/17/07 – TTC emailed Tom Heyl today to tell him that TEI originally signed the payapp on 05/16 but resigned a blank copy on 07/13 at Tony’s office for payment. TEI also asked Process to do some piping work with the remaining $1000 on the PO if Tony says it is ok to do.

01/22/07 – Project complete. TTC not sure final paperwork has been received at this point.

06/27/06 – Bids have started to come in on the job and TTC is tabulated them. TTC will send tabulation to Mike and Tony today when complete.

06/23/06 – TTC sent out prebid meeting notes for project. PDF

06/12/06 – TTC met with Tony and Mike. Prebid on this project is set for June 20. TTC contacted Steve Hutton at GE about attending the prebid. TTC will contact the contractors to schedule the prebid.

by TEI TEI

COURTHOUSE – ROOM 468 JURY COMMISSION OVERFLOW

06/11/07 – TE received the final pay application along with the allowance clarification and close-out paperwork. All materials have been sent to HCFD.

04/27/07 – Mark Donnelly has directed TE to let QCE know that they should send their final pay application so that this project can be closed out through TE. Hamilton County Facilities with deal directly with American Sound to complete the remaining PA system work. TTC reviewing EDi Closeout for submissions to County.

04/13/07 – American Sound will directly quote project to revise PA system wiring to Mark Donnelly of Hamilton County Facilities Department. QCE will be closed out of their involvement with this portion of work.

04/05/07 – QCE is working

03/30/07 – Project reported to have received occupancy from City. Bulletin issued for PA system. Final paperwork from QCE after change order is needed.

03/23/07 – Space is occupied and there a couple of items left to finish and closeout. The PA system has to modified for client satisfaction. TE and MD are working with QCE to resolve the issue. TE and MD agreed on which QCE change order will be paid from the allowance (via email) and what will be a change order to the project (extra conduits).

03/20/07 – Punchlist are complete. Everything is working except the paging system.

03/15/07 – We received nice kudos from Frtizz on this job. Closeout meeting went well. County agreed to insulated the steam piping behind the radiator. TTC and Honeywell agreed that some control tweaking still remains to be done before project is completed. TTC received O&M’s from EDI last Monday and must turn them over after review to MD. Once payapps are received and O&M’s approved this project can be closed out for EDI and QC M.

02/23/07 – TEI distributed the mechanical punchlist for this project to EDI and partners. EDI forwarded to Honeywell and requested everything be completed by 02/28/07 for project completion. PDF

02/15/07 – EDI provided an update on the project as they are trying to do their balance report but the unit has no power yet and the controls are not hooked up. QCE was contacted about having power to the unit and was suppose to do that work yesterday as was Honeywell who did not make it down until today. QCE reported that they did have power to the unit but the fusible disconnect on the 5th needed to be turned on except that no one from QCE was there to turn that on. The balancer is on site today waiting for the system to be turned on to get the readings they need to startup the unit and get the balance report complete. EDI sent in a list of change order and allowance request that TTC processed and forwarded to Mark Donnelly for discussion. This list included the condensate pump, relocation of CHWS/R, steam lines, and fire protection supports. MD and TTC discussed by phone and the County is willing to pay at this point for everything but the relocation of the CHWS/R lines.

01/22/07 – Everyone met onsite today to discuss the MER. Space was allocated for the Controls, Data and Mechanical Systems. Electrical coordinated the locations of the additional power outlet for the mechanical sump pump and the receptacle for the controls. Owner stated that fire support piping will be hung from deck and not form interstitial ceiling in Room 468A. TTC showed EDI this space and EDI will adjust the hangers. Owner will not accept ceiling hung from ductwork. Radiators are complete. In reading room TE asked for Architect to check height of new cover to handle the concrete brick under the legs. Steam system is up and running. 3 of the 4 steam control valves are replaced. One could not be removed because of pipe rupture concerns so TE decided to leave as is. Honeywell submitted a 3-way valve for control on the job but that is not required and it has been plugged to a 2-way valve per TTC’s instruction. The new drain and chilled water piping will be installed above the unit at about seven foot AFF and drop down to the unit next to the unit leaving the unit open for access. We will trade the access door work by the GC for a square opening and two backboards for controls and data in the room. TE coordinated these locations and sizes with CRA and the contractors.

01/17/07 – TEI requested that EDI relocate the CHWS/R away from the disconnects and route over top the unit. QCE and CRA decided that the soffit can be lowered 4″. We will issue bulletins for the additional work and get prices for the County. Honeywell reported that their control panel will only require 8A of service which means it can be fed from our AHU unit or electric heater with a separate transformer and given the low IT requirement, it may share the same circuit as it. TEI called for a coordination mtg on MER on 01/22/07 at 1:00 p.m. to coordinated all these issues.

01/16/07 – TEI is meeting QCE to discuss additional conduits under the soffitt in the CH hallway. TEI invited CRA and HamCo to attend on Tuesday at 4pm.

01/10/07 – Sprinkler contractor was on site working on the sprinkler system on 01/09. HVAC work is progressing. TTC instructed EDI to use the condensate pump provided with the unit, that the integral check valve meets our requirements and to use the 1.25″ piping as called for on the dwgs to drain system. EDI asked if the 3-way valve provided by honeywell could be plugged to be a 2-Way valve and TTC stated that this was okay. EDI reported the the GC would be reworking the ceiling grid to match the dwgs. It is currently running in the wrong direction. The sprinkler contractor have the sprinkler mains ready for final arm overs down to the new ceiling and will come back after ceiling is reworked. EDI inquired today about placing the radiator on concrete blocks, I am not sure this is an idea I am going to approve.

01/02/07 – The AHU was delivered into the MER on 12/29/06. EDI will be working nights to hook up the HVAC to the new unit runs. Honeywell will be in to hook up and download the controls soon but an exact timeline has not been set. TTC asked the County to allow us to use $600 for a steam condensate fix for the old air handlers located in Room 468. MD has given TEI approval to do this work under the $2500 allowance for this project.

12/21/06 – Const Mtg # 2 was held today. EDI has VAV boxes installed, all ductwork hung except for MER. EDI stated FD will be installed when unit is installed. Unit is scheduled to be arrive on December 27. QCE needs to communicate better with field visits

12/14/06 – TEI was onsite today to review mechanical and electrical progress. On the Mechanical side EDI has all the main ductwork installed up to the mechanical room. The VAV boxes are installed and the sprinkler work is installed except for the arm overs that will eventually be connected after the ceiling is installed. The exhaust fan is installed but the new louver is not installed yet but will be next on the agenda. We are waiting for the unit to arrive before more work is completed. We discussed the new radiator location in the small reading room and TEI instructed EDI to try to fit the existing radiator in between the piping and use piping crosses to connect before moving the radiator up or more into the room. They will try this and advise. I do not believe that any controls programming has been installed yet. TEI is meeting with QCE and Mitch on site at 4pm to go over data cabling, power routes and the shutdown of power today for our tie-in for the new AHU unit. Mike Bartlett is keeping an electrician late today to assist in case QCE runs into trouble.

12/04/06 – Work is progressing. EDI has submitted a full schedule which I believe will be incorporated by the GC into the master schedule. EDI informs me that the VAV boxes have been delivered to the site. The unit is still scheduled for 12/28. The project schedule is still valid.

11/21/06 – We held a construction and schedule meeting with the contractors. EDI has began work in the space on the mechanical side. Demo has been completed. The old AHU has been removed along with its ductwork. Holes have been cut into the walls for the new ductwork paths. Equipment is ordered and scheduled to be here mid to late December. Electrically, 3/4″ conduits have been snake across the ceiling including the voice and data for HamCo communications. One conduit in the mechanical opening for the new supply ductwork has to be moved and will be by QCE. QCE cut a hole in the soffit to facilitate new conduit paths. The GC will have to patch this hole at the end of the job. It was discussed and no access panel is needed at this location. TEI stated that VAV box has SCR control on electric heat and this may be done through a PID loop by Honeywell only if they can provide enough stages of heating to keep the space a constant temperature. Fire dampers were discussed with the code official and they are not required but an engineering dwg would be required to change them so it isn’t worth the savings. The FD will be installed.

10/27/06 – Construction has started. Light fixture submittals have been approved. TEI approved final AHU submittals with EDI and Trane. The units size will be decreased by 1 size allowing it to fit into the space but this increased the face velocity through the coil to slight above 500 fpm (a normal max for airflow). Trane assured us that there would be no moisture drift through this unit. All equipment is on order.

10/01/06 – Pre Construction meeting today. EDI turned over submittals. ThermalTech reviewed and sent back comments. Trane AHU is revise and resubmit. EF, Louver, Air Devices are all furnish as corrected. PL is talking to QCE about down sizing to two 3/4″ conduits across the hallway. Mitch requested a conduit for the IT but we also discussed leaving room on the wall in the MER for a patch panel and have one cable go back across the hallway for easier access. The county has already remediated the space so it is ready for work. TTC instructed EDI to have sprinkler piping in MER room ran close to the outside wall to stay away from new electrical disconnect switches in the space. EDI is supposed to find out who the FPC is on this job and let TTC know.

09/01/06 – TTC contacted the City Plan Review to discuss outdoor air calcs. They have reported that we did not use the correct tables for our ventilation tab. Actually I doubled check this and it is the EXACT table he listed. Perhaps they are misconstruing our use of the transient occupancy rule in both the code and the ASHRAE guidelines. This is where I said the owner may have to provide a letter stating that the space is actually only used at full capacity for under 3 hours on any given day – it is OVERFLOW not full use. Bob Martin called me back yesterday and I am calling him back right now to resolve.

06/15/06 – Bids were opened. Mechanical and Electrical are at budget. Once GC bid and is acceptable also. Partners are waiting on a formal approved bid tab to write the contracts for this project

05/25/06 – Drawings arrived today for this project.

05/22/06 – Spencer has submitted the drawings for bidding. Spencer and TTC discussed bidding estimates and the cost were entered and were below the original estimate. QCR called to say that 2 of TEI dwgs were mislabeled as mechanical dwgs when they were really electrical.

04/11/06 – TTC & Sam discussed project design today in office for layout. Sam met with Mike Bartlett yesterday to find a suitable HHW source for reheat. HHW was so far away that it doesn’t make sense for the space design we are using. TTC discussed zoning of the space today with Mike and decided the best and easiest way to serve these spaces based on our learning’s in the 450 Mediation space is to provide two zones with averaging thermostats to keep areas about the same temperature. Ease of use, easy to maintain and both zones can be kept in the MER adjacent to the air handler for ease of maintenance as well. TTC has to confirm with Mark Donnelly and Mike Bartlett economizers for this air handler on Wednesday.

04/11/06 – Final design is being put on the drawings. It appears that electric reheat will be used. TEI will break the room up into 2 zones total. One for mediation and one for the Jury commission. We will use averaging thermostats in the different rooms to adjust the air flow and temperature of the unit serving the space. This is a modular unit with electric reheat on the VAV boxes. A cooling coil and steam coil will provide the space and supply temperature to the main ducts. The two VAV boxes can be located in the new MER we are using for this job.

04/05/06 – TEI has been coordinating with CR and the owner extensively on the requirements of the new space. An owner provided sketch has created a higher density floor plan that Spencer sent to everyone. CR and TEI decided that ductwork can be ran in the space if the ceiling is lowered to 8′-0″. This will allow TEI to run supply and return ductwork in the interstitial 18″ of open space. A modular unit with direct outdoor air can be used to provide VAV boxes with supply air for the space. The radiators can be used for heating in the winter time. Mike pointed out that the reheat system is off during the summer so no reheat is available in the space. TEI will take this into consideration when designing the new system.

03/14/06 – Second Design meeting yielded a more formalized plan. CR will provide a mirrored floor plan (MS suggestion) and TEI will have mechanical and electric work done by April 14th for review with the owner. TEI will most likely replace the lights and the HVAC unit on this project.

03/07/06 – First Design Meeting. Scope defined for space but we need direction from owner on mediation’s part of the space and actual tenants requirements.

02/16/06 – SOSS sent but not approved.

by TEI TEI

CH – JAX Security Camera Upgrade

12/13/07 – TEI received all the closeout paperwork, O&M’s and final pay applications earlier this week. MH, TTC has reviewed and approved. TEI will have these documents sent to Bert Watts for final closeout on this project. TTC sent an email to Copp asking about the remaining project documents for closeout because the Release of Liens, Surety, PW Affidavit were not included.

10/18/07 – TTC received and sent BW certified payroll but no closeout paperwork has been received.

10/04/07 – Terry and Bert still have not received closeout paperwork and payapps for this project. CSI is working on it.

09/03/07 – Bert sent back the final payapp because schedule of values is not itemized. TTC emailed Paul Wilson at Copp and requested new documents and signed allowance usage. TTC sent the allowance usage again by email and filed the final pay application in the HamCo Hold bin.

08/16/07 – TEI sent all final paperwork to Bert. Copp said they would finish and send their final paperwork to Bert.

08/15/07 – TTC reviewed the final payapp and closeout paperwork. Most of the closeout documents look fine. We have not received any O&M’s for approval on this work. I also don’t believe we have a signed allowance usage on this job. We also do not have a signed Punchlist. TTC sent an email to Paul Wilson outlining all the deficiencies on the project closeout scope and told Copp to send the O&M’s directly to Bert Watts. TEI will send the final signed payapp and paperwork to Bert on 08/16.

07/03/07 – TTC contacted Copp systems about closeout paperwork and job forms to use. TEI gave them until July 13 to get everything to TEI for approval.

05/14/07 – Copp submitted a final payapp on this project but it was rejected by ThermalTech for various reasons. TTC emailed Paul the correct document format and closeout paperwork for resubmission.

03/20/07 – Project ready for punchlist. Only training remains.

03/14/07 – Copp stated the the project is complete including our change order work. TTC has to punchlist this project.

02/19/07 – Paul Wilson has the new control panel in. The countertop is being refinished.

02/02/07 – TEI has not heard anything back from the County on the status of this project.

01/22/07 – Sketch was issued in PDF and DWG format to Bert Watts for Dave Spitznagel.

01/10/07 – TTC will issue a sketch based on the field trip on 01/08. TTC also to get prices to add two cameras to day holding cells on 5th floor from Copp. TTC to coordinate new cabinet dimensions with Facility trades.

01/08/07 – Work has progressed to the point where the monitors are installed on the countertop. They are really large and Mark would like to look at changing the CRT monitors to LCD monitors in this project. Dave Spitznagel has the new countertop built that might be skinnier than the old one. A const mtg will be scheduled for Wednesday.

12/28/06 – Paul Wilson called today to inquire about a quad receptacle behind the DVR recorder for this project. He could not find one on the drawings and we all believe there should be one. TTC instructed them and in talking to MWH he agreed that one should be installed. MWH also question why we didn’t spec LCD screens on this job. TTC could not answer as MS did the design but it does seem kind of odd.

12/26/06 – TTC created project journal and filled in back data from previous meetings.

by TEI TEI

CH – SECURE BASEMENT HALLWAYS

10/18/07 -MD sent back final payapps for record on this project with signed change orders.

09/04/07 – The final close out documents where sent to Bert Watts this morning. These items include final pay application, close out documents, O&M manuals with as-built drawings.

08/31/07 – TEI received signed CO#01 from T.J.Williams. Final pay application was reviewed along with all close out documents. All of these items and the O&M manuals will be forwarded to HCFD on Tuesday, September 4.

08/27/07 – TEI sent final change order for signing to TJ Williams. Bert inquired and we told him they were already in the mail last week.

08/22/07 – TJ dropped off as-built dwgs from Copp on the security system work for the project. TEI will integrate into the system and asked for acad dwgs from Copp to do this, not just the hard copies.

08/14/07 – TEI received the modified final pay application along with the signed Change Order No. 01. NR forwarded pay application No. 03 and Change Order No. 01 to Mark Donnelly. TEI still awaiting the closeout documents from T.J. Williams. An email was sent to Diane Vogelsang of T.J. Williams with a list of the final closeout documents on August 8, 2007. NR will contact her again to make sure it was understood that final payment would not be issued until all of these documents were received and forwarded to HCFD.

08/07/07 – TEI received the final 2 pay applications from T.J. Williams. The close-out paperwork was not included with the final pay application. NR will contact them to let them know what is still needed.

08/02/07 – The system was tested and passed last week. T.J. Williams will be submitting the final pay application today. All of these items are being sent to TEI. The only remaining item for TEI is the as-built drawings. T.J. Williams has been told to send the drawings to TEI so that an electronic version of the drawings can be made turned in the HCFD.

07/30/07 – TEI still awaiting word on the installation of the final 2 stair treads.

06/29/07 – TEI still awaiting word on the installation of the final 2 stair treads.

06/22/07 – TEI still awaiting word on the installation of the final 2 stair treads.

06/15/07 – TEI is awaiting word on when the stair treads have been installed so that the system test may be performed. The treads are the only thing holding this project from being closed.

06/08/07 – There are two stair treads that are not installed. When these treads arrive and are installed there will be a system test scheduled. When everything is proven to operated as intended this project can be closed out.

06/01/07 – NR performed a punch list today. The items on the list will be issued today. The stairs on the north side of the building are currently being installed and when they are finished a function check of the security system will be performed by T.J. Williams. Once this is complete and the punch list items are finished the project will be ready to close out.

05/25/07 – TEI will be on site the week of May 28th to complete the punch list.

05/14/07 – Project is complete.

04/27/07 – Issues with the lock hardware has been rectified. Work is progressing as expected.

04/20/07 – Work is progressing as expected.

04/13/07 – Work is progressing as expected. Wiremold issues are still being addressed.

04/06/07 – Work is progressing as expected.

03/30/07 – Work is progressing as expected. Wiremold issues are being addressed.

03/23/07 – Project progressing as expected.

03/13/07 – A construction meeting yielded nice progress. The sprinklers are complete. TJ will use angle to support the new lights in the structure. Bulkhead issues were discussed and Legacy was instructed to have Ken Jones contact Spencer directly for project direction. TTC & PL attended this meeting.

12/04/06 – A pre construction meeting was held. PL is handling the scope on the EE side. TTC will approve the FPC dwgs for the job when they arrive. City requested response from the Architect which has been sent. No submittal data has been received yet.

09/29/06 – Waiting on contracts to be returned from County.

09/22/06 – Contracts have been written and forwarded to county. T.J. Williams was the successful contractor

09/08/06 – Prebid held. Contract is in progress.

09/01/06 – Project was rebid. New prebid was last week. New bids are expected on this job.

by TEI TEI

CH – 6TH FLOOR RENOVATIONS – PHASE II

05/11/08 – TTC faxed and emailed the original approved permit to Bert Watts at the County.

12/20/07 – TEI gave Honeywell a PO to do the FA work on the 6th Floor space. All of this work was not shown on the dwgs and Honeywell prepared a separate PO for the work and gave it to Bert. TE took it from there and will coordinate the installation to make sure it matches our drawings.

11/30/07 – This week we ironed out the Fire Alarm permit and drawing scope with Honeywell for this job. In the future, we (TEI) should be more proactive in getting this work coordinated with our work but I admit that with both Allgeyer and Richmond leaving that I let this little permit issue slip through the cracks. I appreciate Bert and Mike B helping me out on this issue with this project – TTC.

10/17/07 – CRA emailed correction for door hardware access controls code.

10/04/07 – The plumbing permit has been given to the County, the HVAC permit is approved and was emailed to the County, TTC to delivery hard copy on Friday. IBI permit is coming on Friday and then County can call for inspection.

10/01/07 – Building permit approved. CRA delivered the permit to the County Facilities Department.

09/19/07 – TTC sent the HVAC permit to the County and CRA. Don Grace called and said the plumbing permit was approved and given to Bo Blanks. He said there was one small change. TTC contacted 3 FPC firms to move the 4 sprinkler heads in the Phase 2 space. J2 called back and is ready to do the work. TTC gave this information to Mark and Mike to use whomever they wanted. The sprinkler work will not require a permit according to the contractors.

09/10/17 – TTC finished the latest drawings with Friday’s comments. TTC sent them to MD and BW and will print out the permit sets for Tuesday.

08/31/07 – TTC dropped of hard copies of the drawings to Mark Donnelly for his review. TEI is awaiting word back from him on how to proceed from this point.

08/23/07 – TTC sent plumbing fixture cuts to CRA and HamCo for approval. Our floor plans are finished and ready for review. TTC still has a visit to the CH for underslab field work with MB.

08/20/07 – CRA & TEI met with MD to discuss finalizing this project. New project direction was given to the partners. Some of the items covered at this meeting were: Paracubes will be used in the new office area, deep parabolic will not fit. The unisex bathroom has the WC and LAV flipped and TEI can coredrill the floor for the WC and new FD. We will use a floor mounted tank type WC for this job. The refrigerator will need a drip pan that will have its tube ran over to a trap primer connection on the FD. On the HVAC system we discussed removing all the VAV types and just ducting to the existing branches.

08/17/07 – TTC and NR conducted a site visit to review the work area in regards the the changes made to the plans. These changes will be discussed next week.

08/10/07 – TEI will have drawings ready for owner review next Friday, August 17, 2007.

07/17/07 – TTC discussed with MD via email today about MEP completion on this project. TTC agreed to be complete in 3 weeks with the project 2008 project use. CRA sent over a furniture plan for TEI to use for design criteria on this project.

06/29/07 – At the monthly Partnering Meeting with Hamilton County it was decided that the money to perform this work would be budgeted for a 2008 project. TE was directed to complete the design and turn over the drawings to the Hamilton County Facilities Department.

06/15/07 – Electrical field work and design on this project has been completed. Awaiting HVAC work for this project to be issued for owner review.

06/15/07 – NR created project journal and filled in back data.

by TEI TEI

CH – GENERATOR REPLACEMENT PHASE I

02/21/08 – ThermalTech approved the generator submittals from DK on this project today. GM reviewed the Generator submittals and approved them “No Exceptions” 2 copies were mailed back to DK, Wally Gore. 1 copy was sent to Bert Watts and 1 copy was retained by TEI. TEI printed out 6 sets of stamped mechanical dwgs for Triton today and sent them to Grady Reed for permit application on this project.

12/11/07 – We have had a pre construction meeting with Debra Kuempel and Triton on this project

11/30/07 – At last report, this project was in contract approval stage with DK getting the contract on this job as the base bid winning contractor. TEI is waiting to schedule a pre const meeting for this job as soon as the contract is in place and ready to go.

09/10/07 – MD gave TEI the OK to write the contract for Debra Kuempel and Triton Services on only the base bid work for this project. TTC & LKV will write the contract on Tuesday.

09/03/07 – Bids have been opened. MD asked for company portfolios to assist in helping him decide on which contractor to use on this project. TTC called and requested project portfolios form both QCE and DK for the electrical. Triton Services appears to have the lowest price for the mechanical work on this project. TTC sent an excel bid tab to Mark to use to compare total prices on the project.

08/22/07 – TEI finished Addendum 01 to the project which constitutes all the Alternate 01 work on the project. New fuel oil piping system and new fuel oil pumps on the project. In emails, MD has directed us to use manual fill for the belly tank on the job, accept the substitute vendors who may be rejected due to specifications, double contained pipe in the trench, keep the manual fill station for the belly tank. All of this has been incorporated into our Addendum document. Bidder will have to pay and pick up drawings from Queen City Repro for the addendum.

08/08/07 – MD called to instruct TEI to issue and addendum with all the Phase 2 generator work included. TTC brought up quite a few points about this being tough to do at this point. TEI wants to discuss in further detail with Mark. Prebid is next Monday.

07/31/07 – TEI is finalizing the specs and the dwgs today. As with most projects, during the last 10% of the design we find 90% of the info we have needed to do the job all along. TTC has to add generator exhaust insulation to the piping spec and reformat all the PDF files and check the table of contents before sending it to purchasing for bidding. The job is complete. TTC answered MD that the belly tank is 435 gallons, the smallest available. 300 gallons is good for 8 hours so the new tank should last for 11.6 hours.

07/27/07 – We have finished the dwgs and specs and had the TEI QC check performed. After some final changes on Monday, TTC will send to purchasing for bidding.

07/23/07 – TTC requested ITB and SBE numbers from the County. TEI to have final documents ready this Friday for bidding. Advertising and documents should be at Purchasing by Monday July 30th according to TTC email to team on 07/23. TTC & CB to finish mechanical design on this project.

07/16/07 – JP and TA had final review mtg with MD today. New generator is being pushed back until later. For now we are going to install a replacement generator of the same size in the same space. Project is schedule to be completed by next Friday. TTC will assist in the design.

06/30/07 – A review meeting has been set up between TEI and Hamilton County Facilities Department for Monday, July 2nd.

06/15/07 – TEI is awaiting answers from the Building Officials as to what equipment needs to be on the new generator.

06/08/07 – TEI has continued investigating what equipment will be on the new generator.

05/30/07 – TEI coordinating with IBI and building department about who is going to decide if the fire pump has to be put on the new generator. No one is claiming coverage for this piece of the puzzle.

04/27/07 – Loads are being reviewed by TEI. Project is moving along as expected. TTC & TA visited the site and did some additional field work on the space. One concern TTC noted is that the existing Halon system will have to be removed and upgrade if ANY changes are made to the room. According to code officials and inspectors that Halon system must be removed if it is touched. This may cost an additional $10,000 to the project..

04/20/07 – Project is moving along as expected.

04/13/07 – Kickoff meeting was held this week. Project has moved to Design Phase.

04/04/07 – SOSS has arrived signed with a PO. TEI will schedule a kickoff meeting for this project.

03/30/07 – Project still awaiting signed SOSS.

03/23/07 – Project awaiting signed SOSS.

by TEI TEI

CH – ROOM 340/360 HVAC UPGRADE

04/10/08 – Triton submitted closeout paperwork on this project. TTC review and approved. TEi has to write a change order deduct for the allowance money not used on this project and will send that to Triton for processing on this project.

12/28/07 – Grady asked about payapp # 2 and TEI reported that punchlist, O&M’s, payapp 1 have all been received and sent. To closeout this project we need closeout documents, final payapp, certified payroll and an as-built dwg.

12/21/07 – O&M’s have been received by TEI. TTC approved them for distribution to the County. We are sending them to Bert Watts. TTC asked for an additional airflow balancing on the main units themselves for maximum flow with the VFD at 100% to included later.

12/19/07 – JS reported today that the final punchlist work is complete. TEI can now backcheck the punchlist and closeout the job. MB asked about allowance money remaining for moving a piece of equipment at the CH.

12/12/07 – John Sanders faxed over a final signed punchlist to TTC for review yesterday and reported today that the flow switch cut outs have been approved by the code officials. They will now get them installed into the units for final inspection and approval by the city.

11/30/07 – Triton and Honeywell have finished all their punchlist items but the project did not pass final inspection because of the new mechanical code requirement about secondary condensate drainage under cooling coils. This requirement is listed under the General Regulations Section 307 of the OMC for 2007. TTC sent scanned copies of the code to Mike B and Grady R for assistance in getting the permit and finishing this job. Triton reports, “We are getting documentation from carrier as to what they recommend as an overflow protection device to shutdown in the event there is water present outside of the drain pan. The inspector is ok with this as long as carrier is ok with the install – Grady Reed.

10/17/07 – John Sanders called today to tell us that Triton is lowering the piping on AHU-360 and will coordinate that with MB, bringing out a sheet metal cover for the insulation touching the door on CR 360, moving the CHWS label to the correct line and that they have finished the painting of the doors which I told him was fine they way they looked now, . The controls are still being finished

10/08/07 – TTC sent a revised punchlist with picture email to Grady based on his conversation with Mike Bartlett and Willie Schadler. Triton is trying to finish this project but balancing and controls remain to be done.

10/0407 – County said AHU piping was not switched on original air handler. They may mean the piping blocking the filter rack on the second air handler. Regardless, Triton has to change the piping to match the design.

10/01/07 – Triton has stated that have finished punchlist work and controls are finishing their. Payapp # 1 has been received. Grady has the two VFD’s that TEI is supposed to purchase for our error on this job. Units are installed and working very well as far as I know.

09/12/07 – Triton and QCE are on site performing punchlist items. Honeywell has final control parts to install so system can begin running the VFD’s automatically. Randy from Honeywell said he should be able to install the components this week. TTC will check back on progress next week. The CHW piping has not been relocated yet.09/10/07 – TTC inquired about the status of this project with contractors today.

09/03/07 – TTC finished and sent the punchlist to Triton for this project. Controls are not finished and VFD’s are still in manual causing some overcooling of the space. I am confident that the systems will work when complete.

08/22/07 – Both units are now installed in the space. There were some problems on the initial installation but they have now been worked out. The VFD’s were ordered at 480V when the building really has 208V at the units. Triton quick shipped 2 new units that week and the 340 unit was running the end of the week. Air noise and unit noise has been drastically reduced with the new unit but the motor was making a hum that was fixed by changing the tap voltage and reprogramming the VFD. The County asked for new isolation valves in back of the old ones and Triton has provided a change order cost for this. TTC forwarded to County and is awaiting a decision on allowance usage. The CHWS/R piping was connected backwards to the 340 units but Triton flipped it during the 360 changeout work. The 360 changeout work went much better but MB reported that the new CHW piping interferes with filter rack removal. Grady said that Triton will take care of this. MB reported that the motor noise was back and that the judges both complained. Grady had the VFD rep make some changes and thought it was fine. TTC will contact MB to check. TEI is waiting until the controls are finished before punchlisting the job which should be next week. We have coordinated this with MB. TTC told Triton that the dwgs do NOT specifically call out for the steam trap, isolation valve and strainer to be replaced. It was a lineweight issue with Queen City printing that showed them dark.

07/31/07 – TTC recommended the 340 AHU piping be changed during the 360 shutdown weekend work. We can coordinate it so that we do 360 AHU first and if there is time left we can do the 340 AHU. QCE will also change the 30A fuses to 25A during the shutdown. Carrier will investigate the harmonics of the motor noise to get that removed from the unit. AHU seems to be cooling the courtroom just fine.

07/30/07 – NR verified the fuse and wire sizing for QCE. There only modification from the original design is that the disconnect switch frame is larger than intended. The fuse and wire size match the design so the frame size is acceptable.

07/27/07 – After a week of tribulations, the 340 AHU is up and running. There are several issues that have happened on this project that have this changeout tough. TEI spec’d the wrong voltage on the VFD and had the CHWS and CHWR labeled backwards on the unit. Triton did not have all the material they needed. The unit is whisper quiet but the motor makes a funny noise that I think is not right. We are working on resolving these issues and will get this unit checked and balanced before we changeout the 360 unit.

07/23/07 – Triton informed everyone today that TEI spec’d 480V VFD’s while the correct supply voltage to the units is 230V. TTC checked the drawings and we do call for 480V. We don’t know how this happened because we checked this several times during design and coordinated amongst ourselves. Regardless, we have to find 230V VFD’s from Carrier immediately to keep the project moving forward. The old 340 air handler has been removed and Triton was to remove it from site yesterday. The new unit is installed but not powered up yet. Triton is working on the new drive as we speak.

07/18/07 – Triton submitted a $2300 allowance charge for replacing all pneumatic actuators with electric actuators. AKM approved and TTC sent out the paperwork and approved the change. We are scheduled to have a coordination meeting on Friday with Triton for last minute changeout details for the new unit.

06/25/07 – We had our first construction mtg yesterday with our new project manager Grady Reed. John Sanders is going to be our field foreman on this job. We walked the site and Triton is now putting together a schedule and employee list. Mike is checking on best weekends for the judges. TTC is going to distribute the meeting notes for the project.

05/30/07 – Equipment is scheduled to be delivered in mid-June for immediate installation into project.

05/17/05 – TEI finished the control submittals and sent them back to Grady Reed of Triton Services.

03/15/07 – Submittals have been returned to Triton for ordering on this project. Controls asked about 120V power and NR answered via email that 4th floor is fine.

02/21/07 – Pre Const Mtg yesterday. Meeting notes today. PDF

09/01/06 – TTC met Tim Centers on site today to discuss fitting the unit into the space. The only door from the hallway has a maximum width of 35”. The unit must be completely broke down and re-assembled in place. This is possible with the Carrier 39M series although the entire unit must be taken apart. Habegger reported that the unit will be quieter especially since the fan and all associated components are located inside the unit. We are buying a VFD with the unit for speed control. Also we discovered today that the unit when constructed is tall enough so the fan discharge will match up with the existing supply ductwork. There is no need to put a supply elbow on top of the unit. We are also going to keep the existing return and OA as is and butt our unit up against the existing filter rack. There is just enough room for the transition to fit. TTC went over the controls with GD today. The chilled water valve and actuator are new. The steam valve and its actuator may be replaced. The OA damper will remain in use as is.

by TEI TEI

CH – MARBLE STAIR TREAD REPLACEMENT

12/03007 – CRA Punch work completed.

12/05/07 – CRA final pay app and c/o sent to Hummel.

11/14/07 – Awaiting Hummel completion of punch work this week.

10/01/07 – Awaiting Hummel completion of punch work. Awaiting closeout documents. Need As-Built drawings.

09/19/07 – Work completed 9/12. Punchlist issued 09/19.

by TEI TEI

COURTHOUSE Generator Replacement Project

07/29/09 – We have not received final closeout documents from DK on this project yet. A couple of weeks ago, Bert reported that DK was working on closeout and had all the permit requirements to close this project out.

06/22/09 – Triton has sent everything for closeout on this project. DK got all their final permits and approvals, now we just need the remaining closeout documents.

06/16/09 – DK sent in O&M’s and George approved them. We sent them to Bert. TTC called Pam today and DK and said all final payments will be sent to Bert this week. We have not received the Triton closeout docs yet but TTC inquired with Triton about them.

05/22/09 – We are still working on closeout and permits. Jeff Jones approved the configuration and gave it a UL stamp. DK can now get final inspection and complete the project.

02/19/09 – Non Conformance letter from IBI is top issue. GM is working on. Bert wants a separate SOSS to upgrade the building later.

02/02/09 – Final permit sent to Bert Watts last Friday.

01/05/09 – TTC contact DK and Triton about closeout on this project today. DK needs to fix their invoice #3. Triton needs to send in final payapp even though all closeout paperwork is received. TTC sent allowance usage forms to both contractors for $2000 and $5000 use on this project. DK emailed lots of questions about how to fix billing. That is worked out. Bert asked for final inspections to be completed soon and for Triton to get theirs done.

12/20/08 – Bert reports that most of the work is complete. The XFER Switch component is being ordered and special made at the factory.

11/25/08 – DK said work will be starting next week and should wrap up soon.

11/24/08 – TTC contact DK about closeout.

10/16/08 – Debra sent in change order numbers that equal $5000 (the allowance). These have been approved by Bert and TEI. TEI will process the paperwork for usage and send to DK to facilitate the work occurring at the generator project. We will get insulation on the exhaust piping, a refill on the fuel oil and a part replacement on the transfer switch in the room that was not originally part of the scope. TTC dropped of Kohler O&M’s for this project to Bert by hand. Bert called and said he got them.

09/18/08 – O&M’s received and sent to Bert for Triton. Triton is 100% complete with Permit & As-Builts drawings that are being hand delivered to Bert. DK is missing as-builts and permit and need to provide a quote to insulated the exhaust piping.

09/10/08 – County reported issues from load test include loosened bolts, too hot, heat detector tripped and room overheated. Bert wants inspection certificates and we want o insulated the exhaust piping. Triton cannot provide insulation within project budget. BW advised to proceed with DK. DK sent submitted payapp 2. Allowances usage added to pay app page.

08/29/08 – Load bank test is scheduled for 09/06 at 10am. TTC requested all closeout and inspection forms from contractors.

08/25/08 – Vic at Buckeye responded today that they will set up the load bank test for this project directly through Bert.

08/22/08 – Bert wants TEI to call for load bank test. We contact DK via email.

08/08/08 – Both Mechanical and Electrical punchlists have been given to contractors. TEI has finished mechanical as-built dwg.

08/01/08 – TTC sent out the Mechanical Punchlist to Triton and requested closeout paperwork and allowance usage numbers for the upsized generator exhaust. GM sent out the electrical punchlist for this project.

07/29/08 – Generator mini test and training went well. TEI to do punchlist for project. Automatic Transfer switch acts funny but that was not in the scope of this project

07/04/08 – ThermalTech visited the site as Triton & Debra Kuempel installed the equipment. The entire weekend went smoothly.-

by TEI TEI

COURTHOUSE Clerk of Courts Room 315 Renovations

12/31/09 – Archived on 12/31/09 by TTC

04/15/09 – CR Architecture + Design awaiting Trades to complete work so that punchlist/ walk through can occur.

05/05/08 – Issued for permit.

04/21/08 – Preliminary plans sent to Bert Watts for review.

04/07/08 – Approved SOSS received.

03/27/08 – CR Architecture + Design submitted SOSS.

03/10/08 – SOSS requested.

by TEI TEI

CH – Plaza Lighting Study

02/22/10 – Joe Schamer is setting up the const work schedule and will get back to us soon. He is checking with his sub on work.

07/23/10 – TJ finish field work and sent results to TTE

07/28/10 – TTE send report to Bert recommending fix for lighting fixture 6. The only viable option is to replace the conduit between poles 5 & 6.

08/13/10 – Bert asked for paver cost. TEI to contact JB Schmitt for an estimate.

09/29/10 – New dwgs are updated. THP has a set of markups for their work. Should bid out next week.

10/28/10 – Second prebid schedule 10/29 at 11am.

01/24/11 – Pre Const Mtg was held two weeks ago. As soon as the weather warms up, we will begin our work on the plaza lights.

03/30/11- All work has been completed. TTC will check the wiring covering at the Partner meeting.

04/27/11 – TEI is holding final payapps in the hopes of getting the brass covers installed but hasn’t contacted the restorer after passing along the information. We can make this call directly for HamCo to close this out. We need to discuss with Bert extra charges on this project vs. const cost.

06/14/11 – TTC asked SECO if they have anything in warehouse that would match the light post. Have not gotten a response yet. TTC is checking some stores to see if he can find anything. TTC also came up with a new idea, but needs to check something first to see if it will work.

06/29/11 – Seco has not responded from our last communication. We hope to bring our brass cover down on Thursday to review.

07/21/11 – Spoke to local wiremold vendor. Wiremold offers brass raceway, but is a custom product with 250 ft minimum order. Local rep suggested buying standard product and field painting to match.

08/28/11 – We have purchased the materials, now we just need to install the new wiremold.

12/23/11 – Project will be archived for 2012.

by TEI TEI

CH – Room 315 Data Center Unit

10/06/11 – Glen Wilhoit has asked TTC if there is a scope for this project yet. TTC asked Tony M and Bert W what the status of the SOSS is so he can reply to Glen’s question.

12/23/11 – Geiler won the bid. They are working on submitting hard copy bid documents to the county including the performance bond which they state will be over to the County the last week of the year. The contract is waiting to go over to the prosecutors office after final bid documents are received. ThermalTech will accept either horizontal or vertical unit replacement on this project and Geiler is to submit either at the pre-const meeting. There is no allowance on this project at this time because we took the PO out for the bid amount late in the year.

03/27/12 – ThermalTech received submittals from Geiler for their proposed solution. ThermalTech to review the submittals, and issue a Change Order for modifications to the electrical requirements per the submitted equipment.

04/06/12 – Permit was submitted and TTC asked Geiler to call for any inspections needed after the City approves the permit.

04/12/12 – TTC asks Geiler to have their field personnel move the humidifier back to a wall or install it in a better service orientation. Glen at the County said the plenum location their field man installed it behind the unit and he will not be able to service ti in the future. We show it on the wall in a remote mount installation – please get with your man and see if it can be fixed for Hamilton County. Mike Folck can be present to meet with in the field to help coordinate. Geiler sent a suggestion for the problem to ThermalTech. TTC responded that it was acceptable provided – 1.This location does not void the Carrier warranty in any way (mounting equipment on their case). 2. This location does not interfere with inner workings of CRAC unit. 3. This installation meets NEC requirements for electrical. 4. Greg and Glenn give their blessing too.

04/20/12 – The system is up and running and ready for an inspection. ThermalTech will punch this job out on Tuesday (4/24) after the Ameresco meeting if okay with Greg and Tony.

06/20/12 – Close out documents were sent for this project. The remaining items to take care of are: Only 1 O&M was sent, specs require 3 (TE will do the other 2), warranty letter, contact sheet for warranty or service calls, signed punchlist from Geiler, As-built markup.

Payapp #2 is approved and sent to Tony Matre. Payapp #3 is on hold pending the receipt of the above mentioned items.

06/25/12 – All closeout documents received, approved and hand delivered to Tony at Monthly Partner Mtg. ThermalTech completed O&M manuals for Geiler and inserted As-Built, warranty, contact information along with standard equipment care.

09/28/12 – TTC archived this project after the PM meeting.

by TEI TEI

CH – AHU 566 & 580 Replacements

02/22/10 – TTC has to meet with Greg to discuss additional AHU replacements in the Courthouse.

03/30/11 – As soon as the new contract is ready, we will be submitting an SOSS for this work.

05/06/11 – SOSS submitted for approval.

06/29/11 – Kickoff meeting

07/07/11 – We met with Spencer to gather some intel about the existing soffits in the Courtrooms this morning. GBBN knew that these soffits are left over from the old plumbing chases on the 6th floor jail and have always been part of the building. That makes a lot of sense. TEI will believes they have to be incorporated into the new design and has sent diffuser suggestions to Tony for approval. Habegger has been contacted for vendor sizing and budget quotes for the wor

07/22/11 – TTC and MAF completed fieldwork for demo of existing AHU’s and associated ductwork. Also completed fieldwork for layout of new AHU, VAV’s to serve two courtrooms / judge’s chambers on the 5th floor, and routing of hot water piping thru main hallway. Habegger has submitted two options for replacement AHU. Unit will have to be disassembled and rebuilt to fit thru existing doorway on the 4th floor.

08/09/11 – MCB and MAF completed fieldwork for demo of existing electric feeds to AHU’s and fan coil units. Also completed fieldwork for layout of new electric feeds from 5th floor south panel to new AHU, VFD, and VAVs.

08/22/11 – Mech/Elec drawings and specs in the process of TEI internal review.

09/30/11 – Review set given to TTC today for approval. TTC will have ready for bidding and request ITB for public bidding in about 2 weeks.

12/23/11 – Contracts have been approved by HamCo. The project will start in January with a pre-construction meeting.

02/21/12 – TTC has stamped 6 sets of mechanical drawings that will be delivered to Debra-Kuempel for permitting. NV is stamping 3 sets of IBI drawings that will be delivered to QCE. Proceed with permitting this week.

02/23/12 – Lead paint was found in MER. SOV approved for both contractors. Unit will be 80/80 for electric and QCE is okay with this size and feeder wires. No charge.

02/24/12 – DeBra-Kuempel will not disturb during their demo, but SCS construction will have to saw cut the new opening in for the door where the lead paint is located and the electrical and controls contractor may have some conduits, boxes, clamps etc. mounted on the same wall. They were not sure what the industry standard is for removal or remediation. Tony Matre may be able to answer some of the questions as to how the county will proceed. DeBra-Kuempel has everything going through with the background checks, AHU ordered and the permit has been submitted.

03/27/12 – During demo, Queen City found unknown circuits fed from the existing 566 fan coil unit location. These may be feeding added outlets or lights within the judge’s chamber. Depending on the use of these unknown circuits, we have the option to demo, leave in place, or re-route to the appropriate location. Mike B will be investigating further with QCE and update on this issue.

03/28/12 – Mike B met with Tim Ziebold on site. They traced the conduit and it appears to feed an outlet (with nothing plugged in) on the wall in the judges chambers. They were not able to determine if there were any downstream devices connected to this circuit. Thermaltech spoke with Greg Jennings and the plan is as follows: At a time when the courtroom is not in use, and the Judge has left for the day, we shall disconnect the power to the circuit feeding the branch circuit that runs through the disconnect. At that time, we shall go into the chambers and confirm if the circuit feeds the judges computer (via a light on the power strip). If no lights or equipment in the chambers or courtroom are affected, the per Greg Jennings it is OK to leave off at this time. Feed is to re-routed as follows. Intercept feed at corridor wall (north wall of mezz), provide a junction box above door. Route feed east on the north wall to the east wall, turn south and then down to pick up feed where it feeds up, reconnect and turn back on. Any calls or questions should be directed to Mike Barth.04/10/12 – We call for the sprinkler line to remain in the mechanical room on your 566/580 project. After the new ductwork is in and as a matter of fact even with the old ductwork configuration, that one sprinkler might not give you optimal coverage in that room. I spoke to Debra about note #7 on M200 which tells them to work around or modify that sprinkler line under their contract which will leave it in the same place. After they get the ductwork layout done, we may want to address “better” coverage of that space by extending the head up and running a horizontal line over and changing that to two heads that spray down. That was not in their contract to do but I thought we should at least discuss it now. It may come up on final inspection by the building dept if they pay close attention to this one detail. TTC asked Greg who was doing most of the sprinkler repairs. TTC also tells Greg to feel free to direct all engineering type questions to Folck, Barth or himself if he does not know the answer. TTC stopped in on the contractors and answered a few more questions and went over project progress with Clark and Greg (HVAC & Piping).

04/11/12 – The following is a summary of questions from Debra-Kuempel and clarifications regarding this Project:

– New supply/return ductwork to allow for clear access to 5th floor platform and VAV’s through existing doorway. VAV’s to be installed at working height.

– Existing transfer grille about 5th floor driveway is not necessary for current operation. Damper should be closed off and left in place.

– Existing 580 Judge’s Chamber flex return combines with 580 Courtroom return ductwork. No new ductwork needed for 580 Judge’s Chamber return, leave as is.

– New diffuser locations in 580 Courtroom were not spaced evenly in ceiling bulkhead due to existing structural supports, piping and wiring. New diffuser locations are acceptable and were approved in ceiling bulkhead.

– Existing sprinkler line within mechanical room to be worked around or relocated if necessary per note #7, Drawing M200. Existing coverage within mechanical room may not be adequate. Thermal Tech will address final sprinkler coverage with Debra-Kuempel and County when ductwork layout is complete.

– Location of existing chilled water isolation valves are unknown. Debra-Kuempel is tracing chilled water piping to find nearest location valves. Mike Folck attached pdf’s of 5th floor chilled water piping layout. Existing valves are not indicated.

– New hot water supply/return piping to be routed above ceiling south side of hallway. Proposed routing by Debra-Kuempel is acceptable and was approved.

– New air handling unit to be placed along west wall of mechanical room, allowing space for existing/new condensate drains.

– Total external static of air handling unit is 2.96 in wg. Clearance beneath unit is expected to be 4″. This should allow sufficient space for condensate trap without raising unit or removing part of existing floor.

– Thermal Tech/THP will be providing Debra-Kuempel with clarification drawing regarding new 5th floor platform. Debra-Kuempel will provide new railing along back wall openings for fall protection.

– Building permit not approved by historical conservation. Thermal Tech has contacted historical conservation for status update.

– Direct all engineering questions or comments to Thermal Tech. Contact Greg Jennings for all shutdowns and building access.

– Thermal Tech will be setting up a construction meeting for next week (4/16-4/20) once all availability is known.

04/16/12 – Terry spoke to Tony about Const Mtg setup for Tuesday. TTC attend meeting and take notes.

04/25/12 – Permit to relocate the sprinkler in the mechanical room has been approved. S.A. Comunale is the sprinkler company for the courthouse and has been quoted not to exceed $1,400 to relocate it. TTC questioned if the old head could be grandfathered in or if it needed to be changed. TTC offered to call and discuss if necessary.

05/14/12 – TTC sent out the Construction Meeting notes #1, #2, and #3.

05/30/12 – QCE on site to connect external VFD bypass switch, VFD and AHU this week.

06/25/12 – Initial punchlist complete. Inspections underway for final. Sprinkler work scheduled for this Thursday. Smoke detectors and other punchlist items to be addressed soon.

06/28/12 – TEI did walk through and the items that need to be addressed are:

1. Unit is not started. Start unit and provide all startup reports in O&M.

2. Unit is not programmed. Programmed unit to TEI and Hamco Sequences issued in the submittals. Commission unit to meet all published sequences. Demonstrate to owner unit operates as scheduled. In particular that these units operate on local occupancy sensors then by override when needed. Outdoor air is used for economizer only and VAV boxes work to representative zones. VFD on unit programmed and operating according to sequences.

3. Pain door, install hardware, lock door to master key system given by owner.

4. Touch up the condensate drain work. The hub has been cut too low. Instead only use 45 deg ell at outlet or have cond drain PVC pipe 45 into hub. Put standard hub back so water level does not crest over top of lip. Maintain air gap but don’t cut the hub drain down to achieve it.

5. Install fire sprinkler back into room where removed.

6. Conduct all inspections and test. Get final approval permits from IBI and City. Is the smoke detector missing a base or is that the full installed device?

7. Seal the wall above the existing door with fire stop where conduits and piping penetrate.

8. Clean mechanical room.

9. Clean staging area. Remove masonite, clean floors, tough up walls. Restore room to prior to project conditions.

10. Remove walk off mats and door signage.

11. Begin closeout paperwork process.

12. Provide As-Built markups for work and changes.

06/29/12 – Pretest did not happen, technician spent time repairing shorts on the system from demolition of the old systems. He has everything repaired and new parts installed and programmed. Monday the 9th is scheduled to finish checkout of fire alarm system and DDC controls on the air handler. After the test is completed the Inspectors can be scheduled.

07/03/12 – DeBra Kuempel has asked when the final for the smoke alarm test can be set up. Randy Waddell with Honeywell has verified the pretest. Joe has informed them they would like to use the courtroom on Monday 7-9. Randy is completing the controls, and our air balancers will be in on Friday 7-6 to balance. The Fire alarm permit hasn’t been submitted yet for the AHU project with the City. This has always been an easy miss because it is one little blank on the Bldg and Inspections submit App. How can this be rectified prior to Monday? Did Larry say anything about the final test because typically the fire marshal is usually here for those tests and I see on the city permit website we are still in the “not approved” section for that little bit. Reasons for items not approved: The following table lists items not approved along with the most current action taken on the Item. Items may be linked to a history of prior comments. The list of comments can be viewed by clicking on the item like if present (920-Final HVAC Inspection6/19/2012 Partial approval install ok, need test of duct detector, tied to FA system. Queen City spoke with the city building inspector a couple of weeks ago and asked if we would need to have a fire alarm permit. They explained the specific scope of work. He said that all he wanted to see was unit shutdown in a smoke condition. A separate fire alarm permit would not be required. We already have our IBI final. Now that the pre-test has been confirmed, we should be good to go.

07/18/12 – Queen City is working on closeout documents. Punchlist item 1.03 regarding the flexible conduit in question is the control’s contractor’s responsibility. It will be removed by the end of the week. We will have everything over as soon as we can.

07/25/12 – QCE emailed electronic O&M and AS-Builts to ThermalTech on 07/24/12. QCE to forward closeout package ASAP.

09/19/12 – DeBra Kuempel will drop off O&M’s to TEI today. TTC already has the As-Builts and will hold them until Honeywell’s work at the CH has been approved.

09/27/12 – TTC dropped off two 50ft extension cords for Rich to give to Courts to replaced the ones that ended up missing from the construction work.

09/28/12 – TTC discussed closeout with Rich and Tony. The County seems fine with the revised sequences and temperatures. The VFD is running at 20% minimum when occupied and will ramp up when occupancy increases. TTC delivered the O&M documents to Tony Matre by hand on 9/27. This project can be paid out in full.

09/30/12- TTC archived this project on 9/30/12.

by TEI TEI

Courthouse Window Replacement – Consultant to Ameresco

04/25/12 – Project to be archived. THP completed window wind load code study and provided recommendations to GBBN.

3/02/12- Window Removal schedule with HGC to remove existing window in Mechanical Room on Second Floor. Only partially removed but evidence of materials and installation gathered.

03/12/12- Meeting with Ameresco, Facilities and GBBN to discuss status of project.

03/15/12- Meeting with Tim Davis of Aluminum Resources and Beth Sullebarger of The Preservation Group, LLC to discuss historic preservation requirements and scheduling.

03/28/12- Window Mock-up meeting with County to discuss design features of windows and mock-up for presentation to Judges and Building Committee.

by TEI TEI

CH – AHU 566 & 580 Replacements

02/22/10 – TTC has to meet with Greg to discuss additional AHU replacements in the Courthouse.

03/30/11 – As soon as the new contract is ready, we will be submitting an SOSS for this work.

05/06/11 – SOSS submitted for approval.

06/29/11 – Kickoff meeting

07/07/11 – We met with Spencer to gather some intel about the existing soffits in the Courtrooms this morning. GBBN knew that these soffits are left over from the old plumbing chases on the 6th floor jail and have always been part of the building. That makes a lot of sense. TEI will believes they have to be incorporated into the new design and has sent diffuser suggestions to Tony for approval. Habegger has been contacted for vendor sizing and budget quotes for the wor

07/22/11 – TTC and MAF completed fieldwork for demo of existing AHU’s and associated ductwork. Also completed fieldwork for layout of new AHU, VAV’s to serve two courtrooms / judge’s chambers on the 5th floor, and routing of hot water piping thru main hallway. Habegger has submitted two options for replacement AHU. Unit will have to be disassembled and rebuilt to fit thru existing doorway on the 4th floor.

08/09/11 – MCB and MAF completed fieldwork for demo of existing electric feeds to AHU’s and fan coil units. Also completed fieldwork for layout of new electric feeds from 5th floor south panel to new AHU, VFD, and VAVs.

08/22/11 – Mech/Elec drawings and specs in the process of TEI internal review.

09/30/11 – Review set given to TTC today for approval. TTC will have ready for bidding and request ITB for public bidding in about 2 weeks.

12/23/11 – Contracts have been approved by HamCo. The project will start in January with a pre-construction meeting.

02/21/12 – TTC has stamped 6 sets of mechanical drawings that will be delivered to Debra-Kuempel for permitting. NV is stamping 3 sets of IBI drawings that will be delivered to QCE. Proceed with permitting this week.

02/23/12 – Lead paint was found in MER. SOV approved for both contractors. Unit will be 80/80 for electric and QCE is okay with this size and feeder wires. No charge.

02/24/12 – DeBra-Kuempel will not disturb during their demo, but SCS construction will have to saw cut the new opening in for the door where the lead paint is located and the electrical and controls contractor may have some conduits, boxes, clamps etc. mounted on the same wall. They were not sure what the industry standard is for removal or remediation. Tony Matre may be able to answer some of the questions as to how the county will proceed. DeBra-Kuempel has everything going through with the background checks, AHU ordered and the permit has been submitted.

03/27/12 – During demo, Queen City found unknown circuits fed from the existing 566 fan coil unit location. These may be feeding added outlets or lights within the judge’s chamber. Depending on the use of these unknown circuits, we have the option to demo, leave in place, or re-route to the appropriate location. Mike B will be investigating further with QCE and update on this issue.

03/28/12 – Mike B met with Tim Ziebold on site. They traced the conduit and it appears to feed an outlet (with nothing plugged in) on the wall in the judges chambers. They were not able to determine if there were any downstream devices connected to this circuit. Thermaltech spoke with Greg Jennings and the plan is as follows: At a time when the courtroom is not in use, and the Judge has left for the day, we shall disconnect the power to the circuit feeding the branch circuit that runs through the disconnect. At that time, we shall go into the chambers and confirm if the circuit feeds the judges computer (via a light on the power strip). If no lights or equipment in the chambers or courtroom are affected, the per Greg Jennings it is OK to leave off at this time. Feed is to re-routed as follows. Intercept feed at corridor wall (north wall of mezz), provide a junction box above door. Route feed east on the north wall to the east wall, turn south and then down to pick up feed where it feeds up, reconnect and turn back on. Any calls or questions should be directed to Mike Barth.04/10/12 – We call for the sprinkler line to remain in the mechanical room on your 566/580 project. After the new ductwork is in and as a matter of fact even with the old ductwork configuration, that one sprinkler might not give you optimal coverage in that room. I spoke to Debra about note #7 on M200 which tells them to work around or modify that sprinkler line under their contract which will leave it in the same place. After they get the ductwork layout done, we may want to address “better” coverage of that space by extending the head up and running a horizontal line over and changing that to two heads that spray down. That was not in their contract to do but I thought we should at least discuss it now. It may come up on final inspection by the building dept if they pay close attention to this one detail. TTC asked Greg who was doing most of the sprinkler repairs. TTC also tells Greg to feel free to direct all engineering type questions to Folck, Barth or himself if he does not know the answer. TTC stopped in on the contractors and answered a few more questions and went over project progress with Clark and Greg (HVAC & Piping).

04/11/12 – The following is a summary of questions from Debra-Kuempel and clarifications regarding this Project:

– New supply/return ductwork to allow for clear access to 5th floor platform and VAV’s through existing doorway. VAV’s to be installed at working height.

– Existing transfer grille about 5th floor driveway is not necessary for current operation. Damper should be closed off and left in place.

– Existing 580 Judge’s Chamber flex return combines with 580 Courtroom return ductwork. No new ductwork needed for 580 Judge’s Chamber return, leave as is.

– New diffuser locations in 580 Courtroom were not spaced evenly in ceiling bulkhead due to existing structural supports, piping and wiring. New diffuser locations are acceptable and were approved in ceiling bulkhead.

– Existing sprinkler line within mechanical room to be worked around or relocated if necessary per note #7, Drawing M200. Existing coverage within mechanical room may not be adequate. Thermal Tech will address final sprinkler coverage with Debra-Kuempel and County when ductwork layout is complete.

– Location of existing chilled water isolation valves are unknown. Debra-Kuempel is tracing chilled water piping to find nearest location valves. Mike Folck attached pdf’s of 5th floor chilled water piping layout. Existing valves are not indicated.

– New hot water supply/return piping to be routed above ceiling south side of hallway. Proposed routing by Debra-Kuempel is acceptable and was approved.

– New air handling unit to be placed along west wall of mechanical room, allowing space for existing/new condensate drains.

– Total external static of air handling unit is 2.96 in wg. Clearance beneath unit is expected to be 4″. This should allow sufficient space for condensate trap without raising unit or removing part of existing floor.

– Thermal Tech/THP will be providing Debra-Kuempel with clarification drawing regarding new 5th floor platform. Debra-Kuempel will provide new railing along back wall openings for fall protection.

– Building permit not approved by historical conservation. Thermal Tech has contacted historical conservation for status update.

– Direct all engineering questions or comments to Thermal Tech. Contact Greg Jennings for all shutdowns and building access.

– Thermal Tech will be setting up a construction meeting for next week (4/16-4/20) once all availability is known.

04/16/12 – Terry spoke to Tony about Const Mtg setup for Tuesday. TTC attend meeting and take notes.

04/25/12 – Permit to relocate the sprinkler in the mechanical room has been approved. S.A. Comunale is the sprinkler company for the courthouse and has been quoted not to exceed $1,400 to relocate it. TTC questioned if the old head could be grandfathered in or if it needed to be changed. TTC offered to call and discuss if necessary.

05/14/12 – TTC sent out the Construction Meeting notes #1, #2, and #3.

05/30/12 – QCE on site to connect external VFD bypass switch, VFD and AHU this week.

06/25/12 – Initial punchlist complete. Inspections underway for final. Sprinkler work scheduled for this Thursday. Smoke detectors and other punchlist items to be addressed soon.

06/28/12 – TEI did walk through and the items that need to be addressed are:

1. Unit is not started. Start unit and provide all startup reports in O&M.

2. Unit is not programmed. Programmed unit to TEI and Hamco Sequences issued in the submittals. Commission unit to meet all published sequences. Demonstrate to owner unit operates as scheduled. In particular that these units operate on local occupancy sensors then by override when needed. Outdoor air is used for economizer only and VAV boxes work to representative zones. VFD on unit programmed and operating according to sequences.

3. Pain door, install hardware, lock door to master key system given by owner.

4. Touch up the condensate drain work. The hub has been cut too low. Instead only use 45 deg ell at outlet or have cond drain PVC pipe 45 into hub. Put standard hub back so water level does not crest over top of lip. Maintain air gap but don’t cut the hub drain down to achieve it.

5. Install fire sprinkler back into room where removed.

6. Conduct all inspections and test. Get final approval permits from IBI and City. Is the smoke detector missing a base or is that the full installed device?

7. Seal the wall above the existing door with fire stop where conduits and piping penetrate.

8. Clean mechanical room.

9. Clean staging area. Remove masonite, clean floors, tough up walls. Restore room to prior to project conditions.

10. Remove walk off mats and door signage.

11. Begin closeout paperwork process.

12. Provide As-Built markups for work and changes.

06/29/12 – Pretest did not happen, technician spent time repairing shorts on the system from demolition of the old systems. He has everything repaired and new parts installed and programmed. Monday the 9th is scheduled to finish checkout of fire alarm system and DDC controls on the air handler. After the test is completed the Inspectors can be scheduled.

07/03/12 – DeBra Kuempel has asked when the final for the smoke alarm test can be set up. Randy Waddell with Honeywell has verified the pretest. Joe has informed them they would like to use the courtroom on Monday 7-9. Randy is completing the controls, and our air balancers will be in on Friday 7-6 to balance. The Fire alarm permit hasn’t been submitted yet for the AHU project with the City. This has always been an easy miss because it is one little blank on the Bldg and Inspections submit App. How can this be rectified prior to Monday? Did Larry say anything about the final test because typically the fire marshal is usually here for those tests and I see on the city permit website we are still in the “not approved” section for that little bit. Reasons for items not approved: The following table lists items not approved along with the most current action taken on the Item. Items may be linked to a history of prior comments. The list of comments can be viewed by clicking on the item like if present (920-Final HVAC Inspection6/19/2012 Partial approval install ok, need test of duct detector, tied to FA system. Queen City spoke with the city building inspector a couple of weeks ago and asked if we would need to have a fire alarm permit. They explained the specific scope of work. He said that all he wanted to see was unit shutdown in a smoke condition. A separate fire alarm permit would not be required. We already have our IBI final. Now that the pre-test has been confirmed, we should be good to go.

07/18/12 – Queen City is working on closeout documents. Punchlist item 1.03 regarding the flexible conduit in question is the control’s contractor’s responsibility. It will be removed by the end of the week. We will have everything over as soon as we can.

07/25/12 – QCE emailed electronic O&M and AS-Builts to ThermalTech on 07/24/12. QCE to forward closeout package ASAP.

09/19/12 – DeBra Kuempel will drop off O&M’s to TEI today. TTC already has the As-Builts and will hold them until Honeywell’s work at the CH has been approved.

09/27/12 – TTC dropped off two 50ft extension cords for Rich to give to Courts to replaced the ones that ended up missing from the construction work.

09/28/12 – TTC discussed closeout with Rich and Tony. The County seems fine with the revised sequences and temperatures. The VFD is running at 20% minimum when occupied and will ramp up when occupancy increases. TTC delivered the O&M documents to Tony Matre by hand on 9/27. This project can be paid out in full.

by TEI TEI

COURTHOUSE Vestibule Heaters

Partner Lead: TEI

10/21/13 – Fieldwork started for vestibule heaters.

10/31/13 – Review drawings supposed to be delivered to County.

01/30/14 – Tony informed TEI to award just Alternate 1 to Debra. DK sent submittal for approval today. TEI working on contract for DK.

02/25/14 – Matre advised ThermalTech to hold on reviewing submittal until contract was approved. No word on contract details yet.

03/18/14 – Tony, Terry and Dave Gleason discussion moving forward with this project. The contract is at the Prosecutors office for approval, Tony has the option of taking out just a PO to do this work, TTC thought either would be fine for this work.

04/24/14 – Debra has received PO from Tony. TEI to review submittals and get project started with pre-const mtg.

05/28/14 – Submittals received on new equipment.

06/24/14 – Tony says project is a go. TTC to start submittal review for Debra.

07/07/14 – TTC is reviewing the submittal for the center vestibule now.

08/15/14 – Submittal looks good. TTC has to get DK to order the equipment.

09/26/14 – Equipment ordered. Color selection made by Tony, forwarded to DK by TTC. Install schedule in Nov/Dec.

12/09/14 – Debra is on site to install the new vestibule door heater.

12/10/14 – Debra called with an installation question on mounting and heights of the heater. TTC met DK and Tony on site for resolution. The heater will be surface mounted to a new panel DK is providing that matches the existing size of the opening from the old heater. At 7-3/4″ deep this leaves about 16-18″ above people heads when they come through the center vestibule door at the Courthouse. ThermalTech spec’ed this unit with variable outlet temperature and airflow to balance in the field to a flow and temp the owner is comfortable with. The new panel and mounting will be color coded to match but it is a slightly different color then the bronze vestibule metal. DK to come back 12/11/14 to install.

12/11/14 – Debra was called away on an emergency service called. Gleason stated they will be back Friday to install. Electricians to arrive on Monday to pull the additional feeder required for the 3Ph service of the new new heater. The new heater is also much lighter than the old heater, but DK will still provide unistrut support bars across the top of the vestibule to span the weight across on the wall not just the mullions.

01/06/15 – Vestibule heater is installed and working. TTC delivered field report to Tony and Ken about heat output and conditions in lobby. The heater is not on full yet and the lobby is still cold in the dead of winter at the guard shack stand. The new heater does heat up the vestibule in 30 seconds however when the doors are closed.

02/24/15 – Project is complete. Punchlist was issued and TTE confirmed with Ken the heater is on full and working. Ready to archive.

by TEI TEI

COURTHOUSE Bronze Door Cleaning

Partner Lead: GBBN

09/13/13 – Tony Matre requested an SOSS from GBBN to clean the Bronze Doors of the Courthouse.

09/13/13 – GBBN visited the Courthouse to examine the various doors and take photographs. GBBN sends scope questions to Tony Matre for clarification.

09/16/11 – Tony Matre sends answers to GBBN and suggests meeting on site to review the doors, windows and light posts.

09/20/13 – GBBN meets with Tony Matre to view the doors, windows and light posts. Tony presents an earlier vendor quote to GBBN with contact information.

09/25/13 – GBBN completes and delivers SOSS to Bert Watts for processing.

03/27/14 – TEI sent GBBN CH Vestibule Project Manual for coordination on cleaning scope. Only the center vestibule will be done if any are approved. This should not change the cleaning scope in TTC’s opinion.

05/02/14 – GBBN forwards preliminary drawings and specifications to Tony Matre for review.

07/30/14 – Tony Matre calls GBBN to discuss replacement options on hold-open latches for the existing Riot Doors. Tony to send digital photos to GBBN for inquiries with vendors.

by TEI TEI

COURTHOUSE & 230 Site Security Camera Project

Partner Lead: TEI

11/21/14 – SOSS sent to Bert Watts for approval.

12/11/14 – TEI receives approved SOSS and PO.

12/16/14 – TEI meets with Rodney Lofland on site to discuss full scope and deliverables.

01/29/15 – TEI progress meeting and field work with Rodney Lofland and review 80% drawings for project.

02/24/15 – TEI working to complete package. meetings with City, Vicon and integrator scheduled for 2/27 for coordination.

03/25/15 – TEI coordinated with the City of Cincinnati, division of Traffic Safety, and the CPD re the mounting of the cameras on the street lights. City of Cincinnati (Greg Long) told TEI that he does not have authority to approve any cameras on city light poles. TEI talked with Rodney Lofland, who is coordinating with county administrator to see if there are contacts that he has made that can approve the cameras on the light poles. TEI to meet with Rodney on 03/26/15 at 1:00 PM to develop a backup plan if city will not allow on their poles.

04/29/15 – Project was issued for bid on 04/23/15. Pre-Bid conference scheduled for 05/07/15 at the Courthouse. Bid opening date scheduled for 05/21/15 at 11:00 AM at 138 East Court Street.

05/26/15 – Bids received. TEI calling Debra Kuempel (apparent low bid) to qualify bid and drafting up contract for prosecutor approval. Upon return of approved contract, work to be scheduled and installation to begin (assuming around 07/01/15).

06/24/15 – Rodney said the contract is signed. Project commencement waiting on PO from purchasing department. Rodney Lofland indicated about 3 weeks for the turn around of the PO.

07/22/15 – Purchase order signed and sent to DK for project (PO# PO563836). Project pre-construction meeting was conducted on 7/20, and the work shall commence within the next few days according to DK. DK indicated that all new NVRs will be installed and allowed to run with no existing cameras connected to them for a minimum of 1 week (each) to make sure that there are no issues.

08/25/15 – Project is on-going. DK is proceeding with the installation of indoor cameras, NVR’s, Nucleus, and cabling in all of the buildings. Intent is to do all outdoor work (with lift) at same time. Additional cameras were added in Admin Lobby, and (3) 24 port POE network switches were added to project to isolate camera network. Cost for scope changes are allowance usages.

09/23/15 – Project is on-going. DK instructed to have all work (with the exception of skywalk cameras) complete, on line and recording prior to October 1. Punch list site visit scheduled for 10/08/15 with TEI.

10/28/15 – Project installation is complete. All cameras and DVRs are on line and recording. Because of vacation schedules and the DK foreman being on Jury Duty, the punch list has been re-scheduled for 10/29/15 in the afternoon. ThermalTech and Rodney Lofland to perform punchlist and develop write up for DK to complete. Project completion pending closeout documentation.

12/07/15 – Project punch list complete, and TEI is awaiting signed off copy of punch list and closeout documentation for project completion.

01/26/16 – ThermalTech receives as-builts and O&M manuals for review. ThermalTech rejects as-builts – and meets with DK to go over the issues with the drawings. DK working on correcting the as-builts and will re-submit to TEI for approval. TEI to bring to Rodney Lofland upon receipt.

02/23/16 – ThermalTech approved all final closeout documentation and delivers closeout documents, as-built drawings, O&M manuals, signed punchlist and final pay applications to Rodney Lofland. Project completed and ready for closeout.

05/26/16 – Put security cameras on CCTV layers and then archive.

by TEI TEI

COURTHOUSE Law Library Estimating

Partner Lead: GBBN

01/21/16 – Facilities calls GBBN to set up kick-off meeting to discuss project scope. Meeting day, time and location set.

01/25/16 – GBBN met with Mary Jenkins and Bert Watts to review scope and visit spaces to be renovated.

02/22/16 – GBBN delivers SOSS to Facilities.

03/03/16 – GBBN receives approved SOSS and Purchase Order.

03/03/16 – Discussions between GBBN and Facilities to finalize details on MEP issues and sink locations.

03/04/16 – GBBN begins BIM modeling of existing spaces.

03/15/16 – GBBN meets with Interior Project Management to discuss finishes, furniture and interior design concepts.

03/30/16 – GBBN, Facilities and the Law Library staff met to discuss and review the renderings and fly-by video. A CD was left with the staff containing the video and rendering views. Minor revisions to one rendering view are requested. GBBN to send updated rendering view and Floor Plan with View Indicators via email to client.

04/07/16 – GBBN sends final revised renderings, Opinions of Probable Cost and Floor Plan with Rendering Views to Mary Jenkins.

04/08/16 – Facilities inquires about potential construction cost savings for large overall project estimate verses individual. GBBN and TEI confirm overall project would save on mobilization, general conditions costs, design fees and Archibus automatic percentage costs.

04/11/16 – GBBN transmits same potential cost saving information mentioned above to Mary Jenkins upon her request.

05/20/16 – GBBN receives request from Law Library to resend last invoice due to premature closing of original purchase order.

05/23/16 – Law Librarian indicates to GBBN that Board received designs for new renovations well and is considering moving forward.

06/30/16 – submitted as a 2017 project request.

by TEI TEI

COURTHOUSE Wireless Study

Partner Lead: ThermalTech

09/23/15 – ThermalTech Submitted SOSS to Bert Watts – Pending approval.

10/28/15 – ThermalTech receives approved SOSS and PO from Bert Watts. Bert Watts and ThermalTech are awaiting a kickoff meeting with Shannon Supe et al.

01/26/16 – ThermalTech working on this study after the completion and comments from the 800 Broadway study. ThermalTech stated that a review copy to Bert Watts on 02/05/16.

02/23/16 – ThermalTech submits copy of study to Bert Watts for review and comments. ThermalTech to incorporate BW review comments upon reciept and issue.

03/29/16 – ThermalTech to finalize report and study pending final comments from Shannon.

04/27/16 – ThermalTech receives head count information from Shannon, and can finalize report. Report to be completed and turned over to HamCo early/mid May.

05/25/16 – ThermalTech delivers final report to HamCo Facilities for distribution, and sends facilities the report in electronic format.

06/30/16 – submitted as a 2017 project request.

by TEI TEI

COURTHOUSE Facade Assessment

Partner Lead: THP

04/06/15 – County requested SOSS to assess façade and roofs.

04/28/15 – THP developing SOSS.

06/05/15 – THP issued SOSS.

12/08/15 – THP completed roughly 40% of rappelling survey.

01/28/16 – THP completed field survey and working on report.

02/25/16 – THP is working on the report.

04/25/16 – THP complete thermal scans of Courthouse roofs.

04/28/16 – THP anticipates completing the draft report for the County’s review on May 5.

06/19/16 – Draft report complete for THP in house review

by TEI TEI

Courthouse AHU-21 Refurbishment

Building Operations Project

08/06/15 – Temperature probes put in space to monitor temperatures. Temps and space seemed to be tempered well. Temp and humidity within reason. ThermalTech collected original design criteria from Trane on AH-21 and all the other air handlers ordered at the same time from Trane and shared this with the County.

08/12/15 – Preliminary report sent to Tony Matre. Unit is undersized for space at full load and number of people for space. Trace load calc shows 12.5 to 15 tons of cooling required. Existing unit is 5-7.5 tons but space seemed conditioned much better than this.

08/19/15 – Tony requested new SOSS to replace AH-21.

09/23/15 – Tony Matre asked for an SOSS to replace AH-21.

10/14/15 – We decided to bid this project as a Small Project bid.

10/28/15 – ThermalTech met with Facilities to start filed work today on this project for small project bid.

11/25/15 – All bids came in above $50,000. Bidders were advised that no contract could be awarded. ThermalTech contact Trane immediately to inquire about a rebuild project under State Term we had discussed previously. Trane started working on a new proposal immediately.

12/04/15 – Trane submitted a new price. TEI approved scope and cost for a preliminary PO for this work. New coil and motor should help meet max conditions in this room.

01/12/16 – TEI sent state term contract in for approval with Tony and Don Webber.

02/17/16 – Don Webber sent back and email saying he was awaiting comments from County to proceed with work.

02/26/16 – Tony says contract just needs to be signed by Trane and we can kickoff the job. TTC will advise Trane.

03/30/16 – Contract and PO approved, equipment is ordered, waiting for Trane to start.

04/15/16 – Equipment is ordered and will be here in a few weeks. Trane getting background checks in order.

05/18/16 – Trane technicians on site setting up project work. There is discussion about the motor upgrade and VFD size.

06/22/16 – TTC sent out the punchlist and asked Trane for feedback on why the bigger motor was not installed. TEI also recommend that Andy increase the SP of AH-21 to 1.5″.

06/30/16 – Trane received the punchlist and thought they many not have add the motor to make the unit work with the the new coil. They are verifying this. TTC dropped off 5 temp sensors on Tuesday to track temperatures in space. Apparently the worst time is coming out of nigh setback on really hot days. TTC and Tony discussed the next steps after the VFD is up and running at 100%.

07/18/16 – TTC sent VFD data to Trane and asked them to continue with the motor replacement as included in the contract. TTC sent summary field report to Andy and Tony for Room 260 showing that the room is not maintaining temperatures when its over 90F outside and cannot reach 72 on Monday until 4:00pm in the afternoon after hot weekends.

07/21/16 – Don Webber from Trane and his technician will be the Courthouse on Friday to look at installing the motor as part of the contract work for AH-21. My fieldwork shows we are not meeting temperatures on over 90F days and coming out of night setback on the weekends.

07/27/16 – Talked to Tony and Andy, Trane was down and checked the motor and VFD and told Tony that something did not match up. I am waiting for more details from Trane as promised.

08/03/16 – TTC check in with Trane on 8/1, no updates yet.

08/11/16 – Trane sent field reports, TTC looked over the reports and responded requesting a phone conference.

08/31/16 – TTC talked to Andy and request more info from Trane. Running the system at 100% still does not cut it on really hot days and hot weekends coming out of nigh setback

09/08/16 – Closeout meeting scheduled with Trane next Friday 9/16.

09/16/16 – ThermalTech met with Trane and Courthouse to discuss next steps. Trane will resheave the fan, ThermalTech will take the air readings next week.

09/27/16 – ThermalTech was on site tonight getting airflow measurements on this unit. TEI to issue field progress report to Tony and Andy. The basic readings were that AH-21 is delivering at least 2500 cfm so it meets the original design.

11/10/16 – We have contacted Trane to resheave the motor again. The extra duct was traced out by Andy and Terry and fees only the hallways, it can probably be throttled back.

11/16/16 – TTC contacted Don Webber again for a status update on this project. He is out until 11/21.

11/25/16 – Trane said that they were sending at technician down to resheave on 11/17.

12/09/16 – Work is complete by Trane and unit is resheaved. TTC and Andy to verify next spring for cooling load capability.

by TEI TEI

COURTHOUSE – Server Room Study

Partner Lead: ThermalTech

08/12/15 – Bert sent approved SOSS to ThermalTech to update the Archibus cost estimate for this project.

08/16/15 – ThermalTech transferred all three old estimates into one new estimate in Archibus for partners to start working on.

08/21/15 – GBBN coordinated new scope with Bert and ThermalTech for estimate phase of project.

08/24/15 – GBBN completed GC estimate in Archibus. ThermalTech completed Fire Protection estimate in Archibus. GBBN sent updated floor plans to ThermalTech with new door locations and walls to update estimate.

08/25/15 – ThermalTech updated part of the HVAC estimate and finished the plumbing estimate n Archibus.

09/24/15 – ThermalTech meeting with Facilities and CMS on Friday to discuss scope and estimates.

10/14/15 – Previous survey data shared with team. ThermalTech to update report and give to Bert for review.

12/08/15 – ThermalTech has started editing the revised copy of the report. Still waiting on return data from Owner, will follow up with Bert on what to do next. We discussed updating the report and floor plans. Revisit the equipment required, floor plan, cooling units, ups systems and structure.

01/25/16 – ThermalTech will finish this report in the first quarter of 2016.

03/09/16 – Bert provided server room worksheets for ThermalTech to begin working on new server room report.

03/29/16 – Bert provided a Duke Rebate for Server room idea for ThermalTech to investigate.

03/30/16 – TTC tabulated the new data sheets and shared them with Bert. New load is about 40 tons maximum including 25% future growth. We are going to verify server counts from other departments. [PDF]

05/18/16 – Bert to schedule a meeting to discuss funding on this project. Design will begin later this year.

06/27/16 – TTC check a minor change to the load calc sheet based on a formula he updated in the JC Server Room project. TTC and Bert discussed adding data rack and cabling cost to estimate #1827. County is discussing funding availability for this project. Cost are in put and updated.

07/15/16 – TTC schedule design time in third quarter 2016. All cabling and additional charges are in the Archibus estimate except possibly a core switch.

07/27/16 – Archibus estimate updated to include racks, fiber installation, and core switches. Facilities sends Municipal Court Courtroom suggested modifications for adding new offices. Facilities requests Archibus estimate to be updated to include these suggested changes. GBBN to update this final piece of work. We will look at room after partner meeting Thursday.

08/25/16 – Cost estimates updated with additional offices.

09/28/16 – No current updates. All information is into the Archibus system and we are on hold.

10/27/16 – Roof 10 being laid up with new roof membrane. Approximately 75% complete.

10/27/16 – Roof 10 change order work includes new steel posts beneath steel stair to provide bearing to structural roof. Also getting price to truncate four 2″-3″ diameter pipes abandon atop roof slab from old RTU. truncating pipes below structural slab to remove additional curbs/saddles from new roof.

10/27/16 – Roof 6 existing roofing system demolished and temporary roof membrane installed. Contractor may start work laying up new roof as early as next week.

10/27/16 – Subcontractor completed sealant work at terra cotta water table adjacent to Roof 10.

10/27/16 – Subcontractor preparing quote for additional 200 linear feet of sealant work at cracked terra cotta. Pieces adjacent to Roof 10. Work needs to mitigate water intrusion behind new roof membrane.

12/29/16 – Awaiting direction from Owner on this and related projects.

02/22/17 – Awaiting direction from Owner on this and related projects.

by TEI TEI

COURTHOUSE – Fire Alarm System Study

Partner Lead: ThermalTech

04/27/16 – ThermalTech Submitted SOSS to Bert Watts – SOSS was approved and PO issued to ThermalTech.

05/25/16 – TEI is instructed by Building Ops that the full replacement of the system will not be completed at this time and that the balance of the SOSS shall be used for CA assistance with Honeywell replacing the non functioning components only at this time.

08/23/16 – BW indicated at last months meeting that the replacement of the Fire Alarm System will be included in a project request with the Server Room Project when the Server Room Project commences.

12/29/16 – Awaiting direction from Owner on this and related projects.

02/22/17 – Awaiting direction from Owner on this and related projects.

by TEI TEI

COURTHOUSE Facade Exam

Partner Lead: THP

02/24/17 – Hamilton County requested SOSS.

03/02/17 – THP provided Hamilton County with a SOSS.

03/09/17 – County approved SOSS.

04/27/17 – THP working on report.

05/25/17 – THP working on report.

06/29/17 – THP anticipated to issue final report.

07/07/17 – THP Final Façade Assessment Report to County.

07/11/17 – THP issued an opinion of probable cost for City façade repair work.

07/17/17 – THP submitted report to City along with fee.

08/17/17 – THP received an approval letter for the submitted Façade Assessment report. In addition, the letter included a violation notice for work to be performed within the next year. Once the work is complete, a report noting the completion of the work needs to be submitted to the City. The due date for the completion of the work and submission of the letter is August 9, 2018.